### v1

# Convergence order of the quantization error for self-affine measures on Lalley-Gatzouras carpets

Abstract.
Let $E$ be a Lalley-Gatzouras carpet determined by a set of contractive affine mappings $\{f_{ij}\}_{(i,j)\in G}$. We study the asymptotics of quantization error for the self-affine measures $\mu$ on $E$. We prove that the upper and lower quantization coefficient for $\mu$ are both bounded away from zero and infinity in the exact quantization dimension. This significantly generalizes the previous work concerning the quantization for self-affine measures on Bedford-McMullen carpets. The new ingredients lie in the method to bound the quantization error for $\mu$ from below and that to construct auxiliary measures by applying Prohorov’s theorem.
Key words and phrases:
convergence order, quantization error, self-affine measures, Lalley-Gatzouras carpets.
2000 Mathematics Subject Classification:
Primary 28A80, 28A75; Secondary 37B10
1. Introduction
The quantization problem consists in the approximation of a given probability measure by discrete probability measures of finite support in $L_{r}$-metrics. We refer to Graf and Luschgy [10] for rigorous mathematical foundations of quantization theory and [16] for its deep background in information theory and some engineering technology. Applications of quantization theory in numerical integration and mathematical finance have been investigated by Gruber [17], Pagès and his coauthors (cf. [8, 27]). A recent breakthrough by Kesseböhmer et al reveals the close connection between quantization theory and the $L^{q}$-spectra for compactly supported probability measures on $\mathbb{R}^{d}$ [21]. One may see [11, 12, 13, 14, 15, 20, 25, 29, 31, 32] for more results on the quantization for fractal measures.
1.1. Quantization error and its asymptotics
Let $|x|$ denote the Euclidean norm of $x\in\mathbb{R}^{d}$. Given $r\in(0,\infty)$, let $\nu$ be a Borel probability measure on $\mathbb{R}^{d}$ with $\int|x|^{r}d\mu(x)<\infty$. We denote by ${\rm card}(B)$ the cardinality of a set $B$. For every $n\geq 1$, we write $\mathcal{D}_{n}:=\{\alpha\subset\mathbb{R}^{d}:1\leq{\rm card}(\alpha)\leq n\}$. For $x\in\mathbb{R}^{d}$ and a set $A\subset\mathbb{R}^{d}$, let $d_{A}(x):=\inf_{y\in A}d(x,y)$. For a Borel function $g$, let $\|g\|_{r}:=(\int|g|^{r}d\nu)^{1/r}$. The $n$th quantization error for $\nu$ of order $r$ can be defined by
$e_{n,r}(\nu):=\inf_{\alpha\in\mathcal{D}_{n}}\|d_{\alpha}(x)\|_{r}.$
By [10, Lemma 3.4], $e_{n,r}(\nu)$ is equal to the minimum error in the approximation of $\nu$ with discrete probability measures supported at most $n$ points, in the $L_{r}$-metric. We refer to [10] for various interpretations of $e_{n,r}(\nu)$ in different contexts.
One of the main goals in quantization theory is to study the asymptotic property of the quantization error, which can be characterized by the $s$-dimensional ($s>0$) upper and lower quantization coefficient:
$\overline{Q}_{r}^{s}(\nu):=\limsup_{n\to\infty}n^{\frac{r}{s}}e_{n,r}^{r}(\nu),\;\underline{Q}_{r}^{s}(\nu):=\liminf_{n\to\infty}n^{\frac{r}{s}}e_{n,r}^{r}(\nu).$
The upper (lower) quantization dimension $\overline{D}_{r}(\nu)\;(\underline{D}_{r}(\nu))$ for $\nu$ of order $r$, is (typically) the critical point at which the upper (lower) quantization coefficient jumps from infinity to zero (cf. [10, 29]). By [10, Proposition 11.3], we have
$\overline{D}_{r}(\nu)=\limsup_{n\to\infty}\frac{\log n}{-\log e_{n,r}(\nu)};\;\;\underline{D}_{r}(\nu)=\liminf_{n\to\infty}\frac{\log n}{-\log e_{n,r}(\nu)}.$
When $\underline{D}_{r}(\nu)=\overline{D}_{r}(\nu)$, we say that the quantization dimension exists and denote the common value by $D_{r}(\nu)$.
Compared with the upper and lower quantization dimension, we are more concerned about the upper and lower quantization coefficient, because they provide us with exact asymptotic order for the quantization error, when they are both positive and finite.
Assuming the open set condition, Graf and Luschgy established complete results for the asymptotics of the quantization error for self-similar measures on $\mathbb{R}^{d}$ (cf. [11]). Their methods and results have greatly influenced almost all the subsequent study on the quantization for fractal measures.
While it is not easy even to estimate the upper and lower quantization dimension for general probability measures, Kesseböhmer et al [21], identified the upper quantization dimension for an arbitrary compactly supported measure with its Rényi dimension at a critical point and provided several sufficient conditions that guarantee the existence of the quantization dimension. This work, along with Feng-Wang’s results in [7], yields that the quantization dimension exists for the self-affine measures on a large class of planar self-affine sets, including Lalley-Gatzouras carpets. Further, the quantization dimension can be implicitly expressed in terms of Feng-Wang’s formula [7, Theorem 2].
However, the results in [21] do not provide us with useful information on the asymptotic order for the quantization error. In the present paper, we will determine the exact convergence order of the quantization error for the self-affine measures on Lalley-Gatzouras carpets in general.
1.2. Lalley-Gatzouras carpets
Let $m\geq 1$ be an integer. For every $1\leq j\leq m$, let $n_{j}$ be a positive integer. Let $b_{j},d_{j},1\leq j\leq m$, be positive numbers satisfying
  1. (A1)
    $\sum_{j=1}^{m}b_{j}\leq 1;\;d_{m}\leq 1-b_{m}$, and
  2. (A2)
    $b_{j}+d_{j}\leq d_{j+1}$ for all $1\leq j\leq m-1$ if $m\geq 2$.
For $1\leq j\leq m$, let $a_{ij},1\leq i\leq n_{j}$, be positive numbers satisfying
  1. (A3)
    $\sum_{i=1}^{n_{j}}a_{ij}\leq 1,\;\max\limits_{1\leq i\leq n_{j}}a_{ij}<b_{j},a_{n_{j}j}\leq 1-c_{n_{j}j}$, and
  2. (A4)
    $a_{ij}+c_{ij}\leq c_{(i+1)j}\;\;{\rm for}\;1\leq i\leq n_{j}-1$ if $n_{j}\geq 2$.
Let $G:=\{(i,j):\;1\leq i\leq n_{j},1\leq j\leq m\}$ and $G_{y}:=\{1,\ldots,m\}$. We consider
(1.3) $f_{ij}(x,y)=(x,y)\left(\begin{array}[]{cccc}a_{ij}&0\\ 0&b_{j}\\ \end{array}\right)+(c_{ij},d_{j}),\;(x,y)\in\mathbb{R}^{2};\;(i,j)\in G.$
With the assumptions (A1)-(A4), the iterated function system $\{f_{ij}\}_{(i,j)\in G}$ satisfies the open set condition with respect to $U:=(0,1)^{2}$: $f_{ij}(U),(i,j)\in G$, are disjoint subsets of $U$. By [18], there exists a unique non-empty compact set $E$ satisfying
$E=\bigcup_{(i,j)\in G}f_{ij}(E).$
The set $E$ is called the self-affine set determined by $(f_{ij})_{(i,j)\in G}=:\mathcal{I}$. This type of fractals were first introduced and well studied by Lalley and Gatzouras [24], as generalizations of Bedford-McMullen carpets (cf. [2, 26]. The Hausdorff and box-counting dimension for $E$ were determined; necessary and sufficient conditions were given to guarantee the finiteness and positivity of the Hausdorff measure. One may see [24] for more details.
Given a positive probability vector $\mathcal{P}=(p_{ij})_{(i,j)\in G}$, there exists a unique Borel probability measure $\mu$ satisfying
(1.4) $\mu=\sum_{(i,j)\in G}p_{ij}\mu\circ f_{ij}^{-1}.$
We call $\mu$ the self-affine measure associated with $\mathcal{I}$ and $\mathcal{P}$. Let $E_{0}:=[0,1]^{2}$. For $l\geq 1$ and $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\in G^{l}$, we write $f_{\sigma}:=f_{i_{1}j_{1}}\circ\cdots\circ f_{i_{l}j_{l}}$. We will call $f_{\sigma}(E_{0})$ a cylinder of order $l$.
In the past decades, self-affine sets and self-affine measures have attracted great interest of mathematicians (cf. [1, 2, 4, 6, 7, 19, 23, 20, 24, 26, 28]). As was noted in the literature, problems concerning typical self-affine sets and self-affine measures are usually difficult. In [7], D.-J. Feng and Y. Wang determined various dimensions for a class of planar self-affine sets and self-affine measures, and established several computable formulas for such measures. These results were then generalized to more general self-affine measures on the plain by Fraser [9] and were further generalized by Kolossváry [22] to self-affine measures on sponges in $\mathbb{R}^{d}$ by introducing a novel pressure function.
So far, the quantization errors for self-affine measures on Bedford-McMullen carpets have been well studied (cf. [20, 30]). Let $[x]$ denote the largest integer not exceeding $x$ and let $n_{0},m_{0}\in\mathbb{N}$ with $n_{0}\geq m_{0}$. Assume that
$G\subset\{0,1,\ldots,n_{0}-1\}\times\{0,1,\ldots,m_{0}-1\};$
$a_{ij}\equiv\frac{1}{n_{0}}\;((i,j)\in G),\;b_{j}\equiv\frac{1}{m_{0}}\;(j\in G_{y}),$
$c_{ij}\equiv\frac{i}{n_{0}}\;((i,j)\in G),\;d_{j}\equiv\frac{j}{m_{0}}\;(j\in G_{y});$
where $G_{y}$ denotes the projection of $G$ onto the $y$-axis. Then the Lalley-Gatzouras carpet $E$ degenerates to a Bedford-McMullen carpet. Let $\xi=[\frac{\log m_{0}}{\log n_{0}}]$ and $q_{j}:=\sum_{i:(i,j)\in G}p_{ij}$. Kesseböhmer and Zhu proved that the quantization dimension $d_{r}$ for $\mu$ of order $r$ is implicitly given by (cf. [20])
$\bigg{(}\sum_{(i,j)\in G}(p_{ij}m_{0}^{-r})^{\frac{d_{r}}{d_{r}+r}}\bigg{)}^{\xi}\bigg{(}\sum_{j\in G_{y}}(q_{j}m_{0}^{-r})^{\frac{d_{r}}{d_{r}+r}}\bigg{)}^{1-\xi}=1.$
In [30], Zhu further proved that $0<\underline{Q}^{d_{r}}_{r}(\mu)\leq\overline{Q}^{d_{r}}_{r}(\mu)<\infty$ holds, in general.
Thanks to the relatively fine structure of Bedford-McMullen carpets, we considered the auxiliary coding space $\Phi_{\infty}:=G^{\mathbb{N}}\times G_{y}^{\mathbb{N}}$ and constructed a Bernoulli product $W$ as an auxiliary measure. Then the upper and lower quantization coefficient for $\mu$ were well estimated via the measure $W$, by going back and forth between $E$ and $\Phi_{\infty}$. In the Lalley-Gatzouras case, however, neither the linear parts nor the translations of the mappings $f_{ij},(i,j)\in G$, need to be constant. This substantially adds to the complexity of the structure of the carpets and the difficulty in analyzing the quantization error.
Combining [7, Theorem 2] and [20, Theorem 1.11], we obtain that, the quantization dimension exists for an arbitrary self-affine measure on Lalley-Gatzouras carpets; and the quantization dimension can be expressed in terms of the $L^{q}$-spectrum of the projection of $\mu$ onto the $y$-axis. However, such expressions seem to be inconvenient in study of the asymptotic order of the quantization error for $\mu$. We will establish a new expression for the quantization dimension in terms of cylinders of the same order. This expression is more convenient for us to construct suitable auxiliary measures which will be used to estimate the upper and lower quantization coefficient.
1.3. Some notations
For $\sigma=((i_{1},j_{1}),\ldots,(i_{k},j_{k}))\in G^{k}$ and $1\leq h\leq k$, we define $|\sigma|=k$ and $\sigma|_{h}:=((i_{1},j_{1}),\ldots,(i_{h},j_{h}))$. Let $k\geq 1$. Let $\theta$ denote the empty word. For $\sigma\in G^{1}$, we define $\sigma^{-}=\theta$. For every $k\geq 2$ and $\sigma\in G^{k}$, we define $\sigma^{-}:=\sigma|_{k-1}$. Let $G^{*}:=\bigcup_{k\geq 1}G^{k}$. If $\sigma,\omega\in G^{*}\cup G^{\mathbb{N}}$ satisfy $\sigma=\omega|_{|\sigma|}$, we say that $\sigma$ is comparable with $\omega$ and write $\sigma\preceq\omega$. If $\sigma\preceq\omega$ and $\sigma\neq\omega$, we write $\sigma\precneqq\omega$. We call $\sigma,\omega\in G^{*}$ incomparable if neither $\sigma\preceq\omega$ nor $\omega\preceq\sigma$. For the words $\tau$ in $G_{y}^{*}:=\bigcup_{k\geq 1}G_{y}^{k}$, we define $\tau|_{h},\tau^{-}$ and the partial order $\preceq$, in a similar manner.
Let $\omega=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\in G^{l}$ and $\tau=(j_{l+1},\ldots,j_{k})\in G_{y}^{k-l}$. We write
$\omega\ast\tau:=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k});$
$\omega_{y}:=(j_{1},\ldots,j_{l}),\;(\sigma\ast\tau)_{y}=(j_{1},\ldots,j_{k});$
$a_{\omega}:=\prod_{h=1}^{l}a_{i_{h}j_{h}},\;b_{\tau}=\prod_{h=1}^{k}b_{j_{h}};\;a_{\theta}=b_{\theta}:=1.$
For $\sigma=\omega\ast\tau$ with $\omega\in G^{*}$ and $\tau\in G_{y}^{*}$, we write $\omega=:\sigma_{L},\tau=:\sigma_{R}$. Define
$\Psi_{l}:=\big{\{}\sigma=\sigma_{L}\ast\sigma_{R}:b_{\sigma_{y}^{-}}\geq a_{\sigma_{L}}>b_{\sigma_{y}},\sigma_{L}\in G^{l},\sigma_{R}\in G_{y}^{*}\big{\}},\;l\geq 1.$
Let $\Psi^{*}:=\bigcup_{l\geq 1}\Psi_{l}$. Let $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})\in\Psi_{l}$. We define
$A_{L,\sigma}:=c_{i_{1}j_{1}}+\sum_{p=2}^{l}\bigg{(}\prod_{h=1}^{p-1}a_{i_{h}j_{h}}\bigg{)}c_{i_{p}j_{p}},\;A_{R,\sigma}:=A_{L,\sigma}+\prod_{h=1}^{l}a_{i_{h}j_{h}};$
$B_{L,\sigma}:=d_{j_{1}}+\sum_{p=2}^{k}\bigg{(}\prod_{h=1}^{p-1}b_{j_{h}}\bigg{)}d_{j_{p}},\;A_{R,\sigma}:=B_{L,\sigma}+\prod_{h=1}^{k}b_{j_{h}}.$
Then a word $\sigma\in\Psi_{l}$ corresponds to an approximate square $F_{\sigma}$ of order $l$:
$F_{\sigma}:=\big{[}A_{L,\sigma},A_{R,\sigma}\big{]}\times\big{[}B_{L,\sigma},B_{R,\sigma}\big{]}.$
For $\sigma,\omega\in\Psi^{*}$, we write $\sigma\preceq\omega$ if $F_{\sigma}\supset F_{\omega}$. In typical Lalley Gatzouras case, the diameters of approximate squares of the same order can be widely different.
We will also consider the auxiliary coding space $\Phi_{\infty}=G^{\mathbb{N}}\times G_{y}^{\mathbb{N}}$. We define
$\Phi_{l}:=\{\mathcal{L}(\sigma):=\sigma_{L}\times\sigma_{R}:\sigma\in\Psi_{l}\},\;\Phi^{*}:=\bigcup_{l\geq 1}\Phi_{l}.$
We write $\mathcal{L}(\sigma)_{L}:=\sigma_{L}$ and $\mathcal{L}(\sigma)_{R}:=\sigma_{R}$. Since confusion could arise, sometimes we simply write $\sigma$ for $\mathcal{L}(\sigma)$. For $\sigma,\omega\in\Phi^{*}$, we write $\sigma\preceq\omega$, if $\sigma_{L}\preceq\omega_{L}$ and $\sigma_{R}\preceq\omega_{R}$. We call $\sigma,\omega\in\Phi^{*}$ incomparable if neither $\sigma\preceq\omega$ nor $\omega\preceq\sigma$. The difference between $\Psi^{*}$ and $\Phi^{*}$ lies in the fact that they are endowed with different partial orders. For $\sigma\in\Phi^{*}$, we define $\sigma]:=[\sigma_{L}]\times[\sigma_{R}]$, where
$\sigma_{L}]:=\{\rho\in G^{\mathbb{N}}:\sigma_{L}\preceq\rho\},\;[\sigma_{R}]:=\{\tau\in G_{y}^{\mathbb{N}}:\sigma_{R}\preceq\tau\}.$
For $\sigma,\omega\in\Phi^{*}$ with $\sigma\preceq\omega$, we have $\omega]\subset[\sigma]$. We define
$|\sigma_{L}\ast\sigma_{R}|=|\sigma_{L}\times\sigma_{R}|:=|\sigma_{L}|+|\sigma_{R}|,\;(\sigma_{L}\times\sigma_{R})_{y}:=(\sigma_{L}\ast\sigma_{R})_{y}.$
Let $(p_{ij})_{(i,j)\in G}$ be the same as in (1.4). For $j\in G_{y}$, we define $q_{j}:=\sum_{i=1}^{n_{j}}p_{ij}$. For $\sigma_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))$ and $\sigma_{R}:=(j_{l+1},\ldots,j_{k})$, we define
$p_{\sigma_{L}}:=\prod_{h=1}^{l}p_{i_{h}j_{h}};\;q_{\sigma_{R}}:=\prod_{h=l+1}^{k}q_{j_{h}}.$
Remark 1.1.
Let $|B|$ denote the diameter of a set $B\subset\mathbb{R}^{2}$ and $B^{\circ}$ its interior.
  1. (1)
    For every $\sigma\in\Psi^{*}$, we have $a_{\sigma_{L}}<|F_{\sigma}|<\sqrt{2}a_{\sigma_{L}}$.
  2. (2)
    For $\sigma,\omega\in\Psi^{*}$, $F_{\sigma},F_{\omega}$ are either non-overlapping (i.e. $F_{\sigma}^{\circ}\cap F_{\omega}^{\circ}=\emptyset)$, or one of them is a subset of the other.
  3. (3)
    For every $\sigma\in\Psi^{*}$, it is well known that $\mu(F_{\sigma})=p_{\sigma_{L}}q_{\sigma_{R}}$.
  4. (4)
    The approximate squares as defined above are slightly different from those in [24, 26]. Our definition will enable us to cross out the possibility of the awkward situation that $\sigma_{L}\precneqq\omega_{L}$ but $\omega_{R}\precneqq\sigma_{R}$. (cf. Lemma 3.1).
1.4. Statement of the main results
Unlike the Bedford-McMullen carpets, the following cases are possible:
  1. (i)
    for some $\sigma,\omega\in\Psi_{l}$, $|\sigma|\neq|\omega|$, even if $\sigma_{L}=\omega_{L}$; this makes the auxiliary measure in [30]—a Bernoulli product, fail to work. We will construct suitable auxiliary probability measures by applying Prohorov’s theorem.
  2. (ii)
    for some $\sigma,\omega\in\Psi^{*}$ with $\sigma_{y}\precneqq\omega_{y}$, we have $|\sigma_{L}|>|\omega_{L}|$ (cf. Example 2.1). This makes the method in [20, Lemma 2] no longer applicable, in case $\min_{j\in G_{y}}n_{j}=1$ and $\max_{j\in G_{y}}n_{j}\geq 2$. We will present a completely new method to construct pairwise disjoint approximate squares so that the quantization error for $\mu$ can be estimated from below. For this purpose, we will apply some ideas from [22, p. 695].
As the main result of the present paper, we will prove
Theorem 1.2.
Let $(f_{ij})_{(i,j)\in G}$ be as defined in (1.3) and $\mu$ the self-affine measure associated with $(f_{ij})_{(i,j)\in G}$ and a positive probability vector $(p_{ij})_{(i,j)\in G}$. Let $s_{r}$ be the unique positive number satisfying
(1.5) $\lim\limits_{l\to\infty}\frac{1}{l}\log\sum_{\sigma\in\Psi_{l}}(p_{\sigma_{L}}q_{\sigma_{R}}a_{\sigma_{L}}^{r})^{\frac{s_{r}}{s_{r}+r}}=0.$
Then we have $0<\underline{Q}_{r}^{s_{r}}(\mu)\leq\overline{Q}_{r}^{s_{r}}(\mu)<\infty$.
Remark 1.3.
(1) The existence of the limit in (1.5) and that of the unique number $s_{r}$ will be proved in Section 3. (2) As a consequence of Theorem 1.2, we obtain that, $D_{r}(\mu)=s_{r}$ and the $n$th quantization error for $\mu$ is of the same order as $n^{-\frac{1}{s_{r}}}$. This substantially generalizes our previous work in [20, 30].
The remaining part of the paper is organized as follows. In Section 2, we present some basic facts on the approximate squares which will be used frequently; Using these facts, we establish some characterizations for the quantization error. In section 3, we are devoted to the construction of auxiliary measures by applying Prohorov’s theorem. In section 4, we establish estimates for the quantization coefficient via the auxiliary measures and complete the proof for Theorem 1.2.
2. Characterizations of the quantization error
In the following, we always assume that $m\geq 2$, to avoid trivial cases. For two variables $X,Y$ taking values in $(0,\infty)$, we write $X\lesssim Y$ ($X\gtrsim Y$), if there exists some constant $C>0$, such that $X\leq CY$ ($X\geq CY$). We write $X\asymp Y$, if we have both $X\lesssim Y$ and $X\gtrsim Y$.
2.1. Some basic facts
Let us begin with an example showing that when $\sigma_{L},\sigma_{\omega}$ are incomparable and $\sigma_{y}\precneqq\omega_{y}$, it can happen that $|\sigma_{L}|>|\sigma_{\omega}|$.
Example 2.1.
Let $m=3,n_{1}=n_{2}=2,n_{3}=1$. Let
$a_{11}=a_{22}=a_{13}=\frac{1}{9};\;a_{21}=a_{12}=\frac{1}{27};\;b_{1}=b_{2}=b_{3}=\frac{1}{3};$
$c_{11}=c_{13}=0;\;c_{21}=\frac{26}{27};\;c_{12}=\frac{1}{9},\;c_{22}=\frac{4}{27};\;d_{1}=0,\;d_{2}=\frac{1}{3};\;d_{3}=\frac{2}{3}.$
For $x\in\mathbb{R}$ and $p\in\mathbb{N}$, we denote by $\{x\}^{p}$ the word of length $p$ with all entries equal to $x$. We define
$\sigma_{L}=((2,2),(1,1),(1,1),\ldots,(1,1))\in G^{12},\;\sigma_{R}=\{3\}^{13};$
$\sigma_{\omega}=((1,2),(2,1),(2,1),\ldots,(2,1))\in G^{9},\;\omega_{R}=\{1\}^{3}\ast\{3\}^{16}.$
Not that $a_{\sigma_{L}}=3^{-24}$ and $a_{\omega_{L}}=3^{-27}$ and that $b_{1}=b_{2}=b_{3}=3^{-1}$. One can see that $|\sigma|=25<|\omega|=28$ and $\sigma_{y}\precneqq\omega_{y}$, while $|\sigma_{L}|=12>9=|\omega_{L}|$.
For two words $\sigma,\omega\in\Phi^{*}$, with $\sigma_{L},\omega_{L}$ comparable, we have
Lemma 2.2.
Let $\sigma,\omega\in\Psi^{*}$ with $\sigma_{L}\preceq\omega_{L}$. Assume that $\sigma_{y},\omega_{y}$ are comparable. Then we have $\sigma_{y}\preceq\omega_{y}$; in other words, $|\sigma|\leq|\omega|$.
Proof.
Let $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})$. Suppose that $\sigma_{L}\preceq\omega_{L}$ and $|\sigma|>|\omega|$. Then for some $p_{1}\geq 0$ and $p_{2}\geq 1$, we have
$\omega=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),\ldots,(i_{l+p_{1}},j_{l+p_{1}}),j_{l+p_{1}+1},\ldots,j_{k-p_{2}}).$
By the definition of $\Psi^{*}$, we have, $\prod_{h=1}^{k-1}b_{j_{h}}\geq\prod_{h=1}^{l}a_{i_{h}j_{h}}>\prod_{h=1}^{k}b_{j_{h}}$. Hence,
$\prod_{h=1}^{k-p_{2}}b_{j_{h}}\geq\prod_{h=1}^{k-1}b_{j_{h}}\geq\prod_{h=1}^{l}a_{i_{h}j_{h}}\geq\prod_{h=1}^{l+p_{1}}a_{i_{h}j_{h}}.$
This contradicts the fact that $\omega\in\Psi_{l+p_{1}}$. ∎
Remark 2.3.
From Lemma 2.2, one can also see that, for $\sigma,\omega\in\Psi^{*}$, we have, either $F_{\sigma}^{\circ}\cap F_{\omega}^{\circ}=\emptyset$, or $F_{\sigma}\subseteq F_{\omega}$, or $F_{\omega}\subseteq F_{\sigma}$. Indeed, if either $y_{\sigma},y_{\omega}$ are incomparable, or $\sigma_{L},\sigma_{\omega}$ are incomparable, then we clearly have $F_{\sigma}^{\circ}\cap F_{\omega}^{\circ}=\emptyset$; otherwise, by Lemma 2.2, we have $F_{\sigma}\subseteq F_{\omega}$, or $F_{\omega}\subseteq F_{\sigma}$.
For convenience, we write
$\underline{a}:=\min_{(i,j)\in G}a_{ij},\;\overline{a}:=\max_{(i,j)\in G}a_{ij};\;\;\underline{b}:=\min_{j\in G_{y}}b_{j},\;\overline{b}:=\max_{j\in G_{y}}b_{j}.$
We clearly have that $\underline{a}<\overline{b}$, but it is possible that $\overline{a}\geq\underline{b}$.
Lemma 2.4.
Let $A_{1}:=[\frac{\log\underline{a}}{\log\overline{b}}]+1$ and $A_{2}:=[\frac{\log\underline{b}}{\log\overline{a}}]+1$. Assume that $\sigma,\omega\in\Psi^{*}$ and $F_{\omega}\subseteq F_{\sigma}$. Then
  1. (i)
    if $|\omega_{L}|=|\sigma_{L}|+1$, we have $0\leq|\omega|-|\sigma|\leq A_{1}$;
  2. (ii)
    if $|\omega_{L}|-|\sigma_{L}|\geq A_{2}$, then we have, $|\omega|\geq|\sigma|+1$.
Proof.
We write $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})$. Since $F_{\omega}\subseteq F_{\sigma}$, we have, $\sigma_{L}\preceq\omega_{L}$ and $\sigma_{y}\preceq\omega_{y}$. It follows that $|\omega|\geq|\sigma|$.
(i) Assume that $|\omega_{L}|=|\sigma_{L}|+1$. Then for some $1\leq i\leq n_{j_{l+1}}$, we have,
$\omega_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(i,j_{l+1}),j_{l+2},\ldots,j_{k},\ldots,j_{k+h}).$
Suppose that $h>A_{1}$. Then $h-1\geq A_{1}\geq 2$. By the definition of $\Psi_{l}$, we deduce
$\prod_{h=1}^{k+h-1}b_{j_{h}}=\prod_{h=1}^{k}b_{j_{h}}\prod_{h=k+1}^{k+h-1}b_{j_{h}}\leq\prod_{h=1}^{l}a_{i_{h}j_{h}}\overline{b}^{A_{1}}<\prod_{h=1}^{l}a_{i_{h}j_{h}}\underline{a}\leq\prod_{h=1}^{l+1}a_{i_{h}j_{h}}.$
This contradicts the fact that $\omega\in\Psi_{l+1}$. Hence, $h\leq A_{1}$ and $|\omega|\leq k+A_{1}$.
(ii) Assume that $\omega_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(i,j_{l+1}),\ldots,(i_{l+h},j_{l+h}))$ with $h\geq A_{2}$. Again, by the definition of $\Psi_{l}$, we have
$a_{\sigma_{L}}\prod_{h=l+1}^{l+h}a_{i_{h}j_{h}}\leq a_{\sigma_{L}}\overline{a}^{A_{2}}<\prod_{h=1}^{k-1}b_{j_{h}}\underline{b}\leq\prod_{h=1}^{k}b_{j_{h}}.$
It follows that $|\omega|\geq|\sigma|+1$. ∎
Now we give an example to illustrate Lemma 2.4 (ii).
Example 2.5.
In Example 2.1, we redefine $a_{13}=\frac{1}{12},b_{3}=\frac{1}{10},d_{3}=\frac{9}{10}$, and leave all the other parameters unchanged. Let
$\sigma_{L}=((1,1),(1,1),(1,1),\ldots,(1,1))\in G^{9},\;\sigma_{R}=\{1\}^{9}\ast 3;$
$\omega_{L}=((1,1),(1,1),(1,1),\ldots,(1,1))\in G^{10},\;\omega_{R}=\{1\}^{8}\ast 3.$
Note that $a_{\sigma_{L}}=3^{-18}$ and $a_{\omega_{L}}=3^{-20}$. Note that
$b_{\sigma_{y}^{-}}=b_{\omega_{y}^{-}}=3^{-18}\geq a_{\sigma_{L}},a_{\omega_{L}}>b_{\sigma_{y}}=b_{\omega_{y}}=3^{-18}\cdot\frac{1}{10}.$
It follows that $\sigma\in\Psi_{9},\omega\in\Psi_{10}$ and $F_{\omega}\subsetneq F_{\sigma}$, but $|\sigma|=|\omega|=19$.
We end this subsection with the following observation on the length of $\sigma_{R}$ for a word $\sigma\in\Phi^{*}$. We will need it in the characterization for the quantization error.
Lemma 2.6.
Let $A_{3}:=\max\limits_{j\in G_{y}}\max\limits_{1\leq i\leq n_{j}}\frac{a_{ij}}{b_{j}}$ and $A_{4}:=\frac{\log A_{3}}{\log\underline{b}}$. For every $\sigma\in\Psi_{1}$, we have $|\sigma_{R}|\geq 1$; for every $\sigma\in\Psi_{l}$, we have $|\sigma_{R}|\geq A_{4}l$.
Proof.
For $(i,j)\in G$, we have $b_{j}>a_{ij}$. This implies that $|\sigma_{R}|\geq 1$ for every $\sigma\in\Psi_{1}$. Let $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})\in\Psi_{l}$. We have
$\underline{b}^{k-l}\leq\prod_{h=l+1}^{k}b_{j_{h}}=\prod_{h=1}^{k}b_{j_{h}}\prod_{h=1}^{l}b_{j_{h}}^{-1}<\prod_{h=1}^{l}\frac{a_{i_{h}j_{h}}}{b_{j_{h}}}\leq A_{3}^{l}.$
It follows that $|\sigma_{R}|=k-l\geq A_{4}l$. ∎
2.2. Characterizations for the quantization error
We call a finite subset $\Gamma$ of $\Psi^{*}$ a finite anti-chain, if $F_{\sigma},\sigma\in\Gamma$, are pairwise non-overlapping; if, in addition, $E\subset\bigcup_{\sigma\in\Gamma}F_{\sigma}$, then we call $\Gamma$ a finite maximal anti-chain in $\Psi^{*}$. For $A,B\subset\mathbb{R}^{2}$, we define
$d_{h}(A,B):=\inf_{(x,y)\in A,(x^{\prime},y^{\prime})\in B}|x-x^{\prime}|,\;d_{v}(A,B):=\inf_{(x,y)\in A,(x^{\prime},y^{\prime})\in B}|y-y^{\prime}|.$
For every $\sigma\in\Psi_{1}$, we define $\sigma^{\flat}:=\theta$. Next, we assume that $l\geq 2$. For $\sigma\in\Psi_{l}$, we write $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})\in\Psi_{l}$. We have
$\prod_{h=1}^{l-1}a_{i_{h}j_{h}}>a_{i_{l}j_{l}}^{-1}\prod_{h=1}^{k}b_{j_{h}}>\prod_{h=1}^{k}b_{j_{h}}.$
Thus, there exists a unique integer $p\geq 0$, such that
$\prod_{h=1}^{k-p-1}b_{j_{h}}\geq\prod_{h=1}^{l-1}a_{i_{h}j_{h}}>\prod_{h=1}^{k-p}b_{j_{h}}.$
We define $\sigma^{\flat}:=\sigma_{L}^{-}\ast(j_{l},\ldots,j_{k-p})$. One can see that $F_{\sigma}\subset F_{\sigma^{\flat}}$. Write
$\underline{p}:=\min_{(i,j)\in G}p_{ij},\;\overline{p}:=\max_{(i,j)\in G}p_{ij},\;\underline{q}:=\min_{j\in G_{y}}q_{j},\;\overline{q}:=\max_{j\in G_{y}}q_{j};$
$\mathcal{E}_{r}(\theta):=1;\;\mathcal{E}_{r}(\sigma):=\mu(F_{\sigma})a_{\sigma_{L}}^{r};\;\overline{\eta}_{r}:=\overline{p}\;\underline{q}^{-1}\overline{a}^{r};\;\underline{\eta}_{r}:=\underline{p}\;\underline{q}^{A_{1}}\underline{a}^{r}.$
From Remark 1.1 (2) and Lemma 2.4, for every $\sigma\in\Psi^{*}$, we have
(2.1) $\underline{\eta}_{r}\mathcal{E}_{r}(\sigma^{\flat})\leq\mathcal{E}_{r}(\sigma)\leq\overline{\eta}_{r}\mathcal{E}_{r}(\sigma^{\flat})<\mathcal{E}_{r}(\sigma^{\flat}).$
For every $n\geq 1$, (2.1) allows us to define
$\Lambda_{n,r}:=\{\sigma\in\Psi^{*}:\mathcal{E}_{r}(\sigma^{\flat})\geq\underline{\eta}_{r}^{n}>\mathcal{E}_{r}(\sigma)\};\;\varphi_{n,r}:={\rm card}(\Lambda_{n,r}).$
Remark 2.7.
(1) For every $n\geq 1$, $\Lambda_{n,r}$ is a finite maximal anti-chain in $\Psi^{*}$. As we did in [20, Lemma 1], it is easy to show that $\varphi_{n,r}\asymp\varphi_{n+1}$. Moreover, for every $\sigma\in\Lambda_{n,r}$, we have, $\underline{\eta}_{r}^{n+1}\leq\mathcal{E}_{r}(\sigma)<\underline{\eta}_{r}^{n}$. (2) Let $l_{n,r}:=\min_{\sigma\in\Lambda_{n,r}}|\sigma_{L}|$. We have, $\underline{\eta}_{r}^{l_{n,r}}<\underline{\eta}_{r}^{n}$, so $l_{n,r}\geq n$.
Remark 2.8.
For $\sigma,\omega\in\Lambda_{n,r}$, if $\sigma_{L},\omega_{L}$ are comparable, then $\sigma_{y},\omega_{y}$ are incomparable, because otherwise, by Lemma 2.2, we would have $F_{\sigma}\subseteq F_{\omega}$, or $F_{\sigma}\subseteq F_{\omega}$, contradicting the fact that $\Lambda_{n,r}$ is a finite anti-chain.
In order to establish estimates for the quantization error, we need to construct a finite anti-chain $F_{n,r}$ out of $\Lambda_{n,r}$, such that $\mathcal{E}_{r}(\tau)\asymp\underline{\eta}_{r}^{n}$ for every $\tau\in F_{n,r}$, and for every pair of distinct words $\tau^{(1)},\tau^{(2)}$ in $F_{n,r}$, the following holds:
(2.2) $d(F_{\tau^{(1)}},F_{\tau^{(2)}})\gtrsim\max\{|F_{\tau^{(1)}}|,|F_{\tau^{(2)}}|\}.$
In [20, Lemma 2], this was done by selecting, for every $\sigma\in\Lambda_{n,r}$, a word $\sigma^{*}$ with $F_{\sigma^{*}}\subset F_{\sigma}$. Unfortunately, the method in [20] strongly relies on the fact that $a_{ij}\equiv n^{-1}\;((i,j)\in G)$. In fact, that method remains valid under the weaker assumption that $a_{ij}\equiv a_{j}\;(1\leq i\leq n_{j})$ for every $j\in G_{y}$, but it is not applicable to general Lalley-Gatzouras case. We will construct $\mathcal{F}_{n,r}$ in a completely different way. Before we proceed with this construction, let us cope with the extreme case that $n_{j}=1$ for every $j\in G_{y}$, where, $\mathcal{F}_{n,r}$ can be defined in a convenient manner.
Lemma 2.9.
Assume that $n_{j}=1$ for every $j\in G_{y}$. Let $\sigma,\omega\in\Psi^{*}$ with $\sigma_{y},\omega_{y}$ comparable. Then we have, either $F_{\omega}\subseteq F_{\sigma}$, or $F_{\sigma}\subseteq F_{\omega}$. In particular, for every pair $\sigma,\omega$ of distinct words in $\Lambda_{n,r}$, we have that $\sigma_{y},\omega_{y}$ are incomparable
Proof.
By the assumptions that $\sigma_{y}\preceq\omega_{y}$ and $n_{j}=1$ for all $j\in G_{y}$, we know that $\sigma_{L},\omega_{L}$ are comparable. The lemma follows easily from Lemma 2.2. ∎
Remark 2.10.
Let $n>2A_{4}^{-1}A_{2}$. Assume that $n_{j}=1$ for every $j\in G_{y}$. For $\sigma\in\Lambda_{n,r}$, we write $\sigma=((1,j_{1}),\ldots,(l,j_{l}),j_{l+1},\ldots,j_{k})$. From Remark 2.7 (2) and Lemma 2.6, we have, $|\sigma_{R}|>2A_{2}$. By Lemma 2.4 (ii), we may define
$\tau_{\sigma}:=\sigma_{L}\ast((1,j_{l+1}),\ldots,(1,j_{l+1+2A2}))\ast(j_{l+2A_{2}+2},\ldots,j_{k},1,m,\ldots,j_{\overline{k}}).$
Then for $\sigma,\omega\in\Lambda_{n,r}$ with $\sigma\neq\omega$, we have
$d_{v}(F_{\tau_{\sigma}},F_{\tau_{\omega}})\geq\underline{b}^{2}\max\{b_{\sigma_{y}},b_{\omega_{y}}\}\geq\underline{b}^{2}(1+\underline{b}^{-2})^{-\frac{1}{2}}\max\{|F_{\tau_{\sigma}}|,|F_{\tau_{\omega}}|\}.$
Thus, in case that $n_{j}=1$ for every $j\in G_{y}$, it is sufficient to define
$\mathcal{F}_{n,r}:=\{\tau_{\sigma}:\sigma\in\Lambda_{n,r}\}.$
Next, we assume that $n_{j_{0}}\geq 2$ for some $j_{0}\in G_{y}$. We will construct $F_{n,r}$ in two steps. For the choice of $\overline{\sigma}$ in the first step, we are inspired by [22, p. 695].
Step 1: For every $\sigma\in\Lambda_{n,r}$, we insert a word $\tau_{0}:=((1,j_{0}),(n_{j_{0}},j_{0}))$ immediately after $\sigma_{L}$. That is, for $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k})$, we define
(2.3) $\overline{\sigma}:=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(1,j_{0}),(n_{j_{0}},j_{0}),j_{l+1},\ldots,j_{k},\ldots,j_{\overline{k}}).$
We define $\mathcal{B}_{n,r}:=\{\overline{\sigma}:\sigma\in\Lambda_{n,r}\}$. We will prove that, for all large $n$, for every pair of distinct words $\overline{\sigma},\overline{\omega}\in\mathcal{B}_{n,r}$, we have, either (2.2) holds with $\overline{\sigma},\overline{\omega}$ in place of $\tau^{(1)},\tau^{(2)}$ (cf. Lemma 2.11), or $\overline{\sigma}^{*}_{y},\overline{\omega}^{*}_{y}$ are incomparable (cf. Lemma 2.14).
Step 2: For every $\overline{\sigma}\in\mathcal{B}_{n,r}$, we select a word $\overline{\sigma}^{*}$ such that $F_{\overline{\sigma}^{*}}\subset F_{\overline{\sigma}}$, and for every pair $\overline{\sigma}^{*},\overline{\sigma}^{*}$ of distinct words, (2.2) holds with $\widetilde{\sigma}^{*},\widetilde{\omega}^{*}$ in place of $\tau^{(1)},\tau^{(2)}$. Then we define $\mathcal{F}_{n,r}:=\{\overline{\sigma}^{*}:\overline{\sigma}\in\mathcal{B}^{\flat}_{n,r}\}$.
Lemma 2.11.
Let $\sigma,\omega\in\Lambda_{n,r}$ with $\sigma_{L},\omega_{L}$ incomparable and $\overline{\sigma}_{y}\preceq\overline{\omega}_{y}$. We have
$d_{h}(F_{\overline{\sigma}},F_{\overline{\omega}})\geq 2^{-\frac{1}{2}}\underline{a}^{2}\max\{|F_{\overline{\sigma}}|,|F_{\overline{\omega}}|\}.$
Proof.
As $\sigma_{L},\omega_{L}$ are incomparable, we have, $f_{\sigma_{L}}(E_{0}),f_{\omega_{L}}(E_{0})$ are non-overlapping. Note that $F_{\overline{\sigma}}\subset f_{\sigma_{L}\ast\tau_{0}}(E_{0})$ and $F_{\overline{\omega}}\subset f_{\omega_{L}\ast\tau_{0}}(E_{0})$. Using this and the assumption that $\overline{\sigma}_{y}\preceq\overline{\omega}_{y}$, we deduce
$d_{h}(F_{\overline{\sigma}},F_{\overline{\omega}})$ $\geq$ $d_{h}(f_{\sigma_{L}\ast\tau_{0}}(E_{0}),f_{\omega_{L}\ast\tau_{0}}(E_{0}))$
$\geq$ $\underline{a}^{2}\max\{a_{\sigma_{L}},a_{\omega_{L}}\}\geq 2^{-\frac{1}{2}}\underline{a}^{2}\max\{|F_{\overline{\sigma}}|,|F_{\overline{\omega}}|\}.$
This completes the proof of the lemma. ∎
In the following we are going to examine the comparability between $\overline{\sigma}_{y},\overline{\omega}_{y}$, when $\sigma_{L},\omega_{L}$ are comparable. We begin with the simplest cases when $|\omega_{L}|-|\sigma_{L}|\leq 2$. By Lemma 2.6 and Remark 2.7 (2), for every $n\geq 4A_{4}^{-1}$ and every $\tau\in\Lambda_{n,r}$, we have $|\tau_{R}|\geq 4$. We will assume that $n\geq 4A_{4}^{-1}$ in the subsequent Lemma 2.12 and Example 2.13 in order to avoid some trivial cases.
Lemma 2.12.
Let $\sigma,\omega\in\Lambda_{n,r}$. Assume that $\sigma_{L}\preceq\omega_{L}$ and $|\omega_{L}|-|\sigma_{L}|\leq 2$, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.
Proof.
By Remark 2.8, when $\sigma_{L}\preceq\omega_{L}$, we have $\sigma_{y},\omega_{y}$ are incomparable. We assume that $|\sigma|=k_{1}$ and $|\omega|=k_{2}$ and write (for $0\leq h\leq 2$)
(2.4) $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),j_{l+1},\ldots,j_{k_{1}});$
(2.5) $\omega=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),\ldots,(\hat{i}_{l+h},\hat{j}_{l+h}),\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{2}}).$
(i) First we assume that $\sigma_{L}=\omega_{L}$, then $\sigma_{R},\omega_{R}$ are incomparable. Note that
$\overline{\sigma}_{y}|_{k_{1}+2}=(\sigma_{L})_{y}\ast(j_{0},j_{0})\ast\sigma_{R},\;\overline{\omega}_{y}|_{k_{2}+2}=(\sigma_{L})_{y}\ast(j_{0},j_{0})\ast\omega_{R}.$
It follows that $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.
(ii) Now we assume that $|\omega_{L}|=|\sigma_{L}|+1$. We write
$\overline{\sigma}:=\sigma_{L}\ast\tau_{0}\ast(j_{l+1},j_{l+2},\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}});$
$\overline{\omega}:=\sigma_{L}\ast(\hat{i}_{l+1},\hat{j}_{l+1})\ast\tau_{0}\ast(\hat{j}_{l+2},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
For convenience, we write $\overline{\sigma}_{y}$ and $\overline{\omega}_{y}$ in detail:
$\overline{\sigma}_{y}=(j_{1},\ldots,j_{l},j_{0},j_{0},j_{l+1},j_{l+2},\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}}),$
$\overline{\omega}_{y}=(j_{1},\ldots,j_{l},\hat{j}_{l+1},j_{0},j_{0},\hat{j}_{l+2},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
If $\hat{j}_{l+1}\neq j_{0}$, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable. Next, we assume that $\hat{j}_{l+1}=j_{0}$. If $j_{l+1}\neq j_{0}$, we again have that $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable; otherwise, we have, $\hat{j}_{l+1}=j_{l+1}=j_{0}$. Note that $\sigma_{y},\omega_{y}$ are incomparable. We deduce that $(j_{l+2},\ldots,j_{k_{1}})$ and $(\hat{j}_{l+2},\ldots,\hat{j}_{k_{2}})$ are incomparable. Hence, $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.
(iii) Finally, we assume that $|\omega_{L}|=|\sigma_{L}|+2$. We write
$\overline{\sigma}:=\sigma_{L}\ast\tau_{0}\ast(j_{l+1},j_{l+2}\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}});$
$\overline{\omega}:=\sigma_{L}\ast((\hat{i}_{l+1},\hat{j}_{l+1}),(\hat{i}_{l+2},\hat{j}_{l+2}))\ast\tau_{0}\ast(\hat{j}_{l+3},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
Then $\overline{\sigma}_{y}$ and $\overline{\omega}_{y}$ take the following form:
$\overline{\sigma}_{y}=(j_{1},\ldots,j_{l},j_{0},j_{0},j_{l+1},j_{l+2},j_{l+3},\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}}),$
$\overline{\omega}_{y}=(j_{1},\ldots,j_{l},\hat{j}_{l+1},\hat{j}_{l+2},j_{0},j_{0},\hat{j}_{l+3},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
If $(\hat{j}_{l+1},\hat{j}_{l+2})\neq(j_{0},j_{0})$, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable. In the following, we assume that $(\hat{j}_{l+1},\hat{j}_{l+2})=(j_{0},j_{0})$. If $(j_{l+1},j_{l+2})\neq(j_{0},j_{0})$, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable. Finally, we assume that $(\hat{j}_{l+1},\hat{j}_{l+2})=(j_{l+1},j_{l+2})=(j_{0},j_{0})$. Because $\sigma_{y},\omega_{y}$ are incomparable, we deduce that $(j_{l+3},\ldots,j_{k_{1}})$ and $(\hat{j}_{l+3},\ldots,\hat{j}_{k_{2}})$ are incomparable. This implies that $\overline{\sigma}_{y},\overline{\omega}_{y}$ are again incomparable, hence the lemma follows. ∎
As the following example shows, things are getting more complicated, when $\sigma_{L},\omega_{L}$ are comparable and $\big{|}|\sigma_{L}|-|\omega_{L}|\big{|}\geq 3$.
Example 2.13.
We assume that $\sigma_{L}\precneqq\omega_{L}$ and $|\omega_{L}|=|\sigma_{L}|+3$. At this moment we temporally do not require that $\sigma,\omega\in\Lambda_{n,r}$. Next, we show that, it is possible that $\sigma_{y},\omega_{y}$ are incomparable, but $\overline{\sigma}_{y},\overline{\omega}_{y}$ are comparable. We write
$\sigma=\sigma_{L}\ast(j_{l+1},j_{0},j_{0},j_{l+4},j_{l+5},\ldots,j_{k_{1}});$
$\omega=\sigma_{L}\ast((\hat{i}_{l+1},j_{0}),(\hat{i}_{l+2},j_{0}),(\hat{i}_{l+3},\hat{j}_{l+3}))\ast(\hat{j}_{l+4},\ldots,\hat{j}_{k_{2}}).$
Here, we have assumed that $j_{l+2}=j_{l+3}=j_{0}=\hat{j}_{l+1}=\hat{j}_{l+2}$. We have
$\sigma_{y}=(j_{1},\ldots,j_{l},j_{l+1},j_{0},j_{0},j_{l+4},\ldots,j_{k_{1}}),$
$\omega_{y}=(j_{1},\ldots,j_{l},j_{0},j_{0},\hat{j}_{l+3},\hat{j}_{l+4},\ldots,\hat{j}_{k_{2}}).$
When $j_{l+1}\neq j_{0}$, one can see that $\sigma_{y},\omega_{y}$ are incomparable, but
$\overline{\sigma}_{y}=(j_{1},\ldots,j_{l},j_{0},j_{0},j_{l+1},j_{0},j_{0},j_{l+4},\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}}),$
$\overline{\omega}_{y}=(j_{1},\ldots,j_{l},j_{0},j_{0},\hat{j}_{l+3},j_{0},j_{0},\hat{j}_{l+4},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
If $\hat{j}_{l+3}=j_{l+1}$, and the two words $(j_{l+4},\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}}),(\hat{j}_{l+4},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}})$ are comparable, then $\overline{\sigma}_{y},\overline{\omega}_{y}$ are comparable.
For the words $\sigma,\omega$ in Example 2.13, we will show in the next lemma that $k_{1}\leq k_{2}$. It is helpful to note that $\omega_{y}|_{k_{1}}$ is a permutation of $\sigma_{y}$. This will actually cross out the possibility that $\sigma,\omega$ belong to $\Lambda_{n,r}$ simultaneously, when $n$ is sufficiently large. Next, we will prove by contradiction that, for sufficiently large $n$ and for every pair of words $\sigma,\omega\in\Lambda_{n,r}$ with $\sigma_{L},\omega_{L}$ comparable, $\overline{\sigma}_{y},\overline{\omega}_{y}$ are necessarily incomparable.
Lemma 2.14.
Let $A_{5,r}:=\big{[}\frac{\log(\underline{q}^{2}\underline{\eta}_{r})}{r\log\overline{a}}\big{]}$ and $n\geq T_{1,r}:=[A_{4}^{-1}(A_{1}+A_{5,r}+3)]+1$. Then for $\sigma,\omega\in\Lambda_{n,r}$ with $\sigma_{L}\precneqq\omega_{L}$, we have that $\overline{\sigma}_{y},\overline{\omega}_{y}$ are incomparable.
Proof.
Assume that $\sigma,\omega\in\Lambda_{n,r},\sigma_{L}\precneqq\omega_{L}$, but $\overline{\sigma}_{y},\overline{\omega}_{y}$ are comparable. We will deduce a contradiction. Let $\sigma,\omega$ be the same as in (2.4) and (2.5) with $h\geq 3$ (cf. Lemma 2.12). Then
$\overline{\sigma}:=\sigma_{L}\ast\tau_{0}\ast(j_{l+1},\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}});$
$\overline{\omega}:=\sigma_{L}\ast((\hat{i}_{l+1},\hat{j}_{l+1}),\hat{i}_{l+2},\hat{j}_{l+2})\ldots,(\hat{i}_{l+h},\hat{j}_{l+h}))\ast\tau_{0}\ast(\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
Claim 1: we have $k_{2}\geq k_{1}$. Suppose that $k_{2}<k_{1}$. Then $k_{1}>l+h$ and
$\overline{\sigma}_{y}=(j_{1},\ldots,j_{l},j_{0},j_{0},j_{l+1},\ldots,j_{l+h-2},j_{l+h-1},j_{l+h},j_{l+h+1}\ldots,j_{k_{1}},\ldots,j_{\overline{k}_{1}}),$
$\overline{\omega}_{y}=(j_{1},\ldots,j_{l},\hat{j}_{l+1},\hat{j}_{l+2},\hat{j}_{l+3},\hat{j}_{l+4},\ldots,\hat{j}_{l+h},j_{0},j_{0},\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{2}},\ldots,j_{\overline{k}_{2}}).$
By the assumption, we have that $\overline{\sigma}_{y},\overline{\omega}_{y}$ are comparable. Hence,
(2.6) $(\hat{j}_{l+1},\hat{j}_{l+2})=(j_{0},j_{0})=(j_{l+h-1},j_{l+h});$
(2.7) $(j_{l+1},\ldots,j_{l+h-2})=(\hat{j}_{l+3},\ldots,\hat{j}_{l+h});$
$(\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{2}})\precneqq(j_{l+h+1},\ldots,j_{k_{1}}).$
Note that $\sigma\in\Psi_{l}$. we have $\prod_{h=1}^{k_{1}-1}b_{j_{h}}\geq a_{\sigma_{L}}>\prod_{h=1}^{k_{1}}b_{j_{h}}$. Hence,
$b_{\omega_{y}}$ $=$ $\prod_{h=1}^{l}b_{j_{h}}\prod_{h=l+1}^{l+h}b_{\hat{j}_{h}}\prod_{h=l+h+1}^{k_{2}}b_{\hat{j}_{h}}$
$=$ $\prod_{h=1}^{l}b_{j_{h}}\bigg{(}b_{j_{0}}^{2}\prod_{h=l+3}^{l+h}b_{\hat{j}_{h}}\bigg{)}\prod_{h=l+h+1}^{k_{2}}b_{\hat{j}_{h}}$
$=$ $\prod_{h=1}^{l}b_{j_{h}}\bigg{(}\prod_{h=l+1}^{l+h-2}b_{j_{h}}b_{j_{0}}^{2}\bigg{)}\prod_{h=l+h+1}^{k_{2}}b_{j_{h}}$
$\geq$ $\prod_{h=1}^{k_{1}-1}b_{j_{h}}>a_{\sigma_{L}}>a_{\omega_{L}}.$
This contradicts the fact that $\omega\in\Psi_{l+h}$ and Claim 1 follows. Thus,
(2.8) $(\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{1}})\preceq(j_{l+h+1},\ldots,j_{k_{2}})\;\;{\rm if}\;k_{1}>l+h.$
Claim 2: we have $\mathcal{E}_{r}(\omega)\leq\underline{q}^{-2}\overline{a}^{hr}\mathcal{E}_{r}(\sigma)$. We distinguish two cases.
Case 1: $k_{1}>l+h$. In this case, using (2.6)-(2.8), we deduce
$\mathcal{E}_{r}(\omega)$ $=$ $\big{(}p_{\sigma_{L}}\prod_{h=l+1}^{l+h}p_{\hat{i}_{h}\hat{j}_{h}}\prod_{h=l+h+1}^{k_{2}}q_{\hat{j}_{h}}\big{)}(a_{\sigma_{L}}^{r}\prod_{h=l+1}^{l+h}a_{\hat{i}_{h}\hat{j}_{h}}^{r})$
$=$ $\big{(}p_{\sigma_{L}}p_{\hat{i}_{l+1}\hat{j}_{l+1}}p_{\hat{i}_{l+2}\hat{j}_{l+2}}\prod_{h=l+1}^{l+h-2}p_{i_{h}j_{h}}\prod_{h=l+h+1}^{k_{2}}q_{\hat{j}_{h}}\big{)}(a_{\sigma_{L}}^{r}\prod_{h=l+1}^{l+h}a_{\hat{i}_{h}\hat{j}_{h}}^{r})$
$\leq$ $\big{(}p_{\sigma_{L}}\prod_{h=l+1}^{l+h-2}q_{j_{h}}\prod_{h=l+h+1}^{k_{1}}q_{j_{h}}\big{)}(a_{\sigma_{L}}^{r}\prod_{h=l+1}^{l+h}a_{\hat{i}_{h}\hat{j}_{h}}^{r})$
$\leq$ $\underline{q}^{-2}\overline{a}^{hr}\mathcal{E}_{r}(\sigma).$
Case 2: $k_{1}\leq l+h$. In this case, we have
$\mathcal{E}_{r}(\omega)$ $\leq$ $\big{(}p_{\sigma_{L}}\prod_{h=l+1}^{l+h-2}q_{j_{h}}\big{)}(a_{\sigma_{L}}^{r}\prod_{h=l+1}^{l+h}a_{\hat{i}_{h}\hat{j}_{h}}^{r})$
$\leq$ $\big{(}p_{\sigma_{L}}\prod_{h=l+1}^{k_{1}-2}q_{j_{h}}\big{)}(a_{\sigma_{L}}^{r}\prod_{h=l+1}^{l+h}a_{\hat{i}_{h}\hat{j}_{h}}^{r})$
$\leq$ $\underline{q}^{-2}\overline{a}^{hr}\mathcal{E}_{r}(\sigma).$
Claim 3: we have $h\leq A_{5,r}$. Assume that $h>A_{5,r}$. Then by Claim 2, we have $\mathcal{E}_{r}(\omega)<\underline{\eta}_{r}^{n+1}$, a contradiction (cf. Remark 2.7).
Using Claims 1-3, we are able to complete the proof of the lemma. We again distinguish between two cases.
Case (i): $A_{5,r}<3$. In this case, we have, $3\leq h\leq A_{5,r}<3$, a contradiction.
Case (ii): $A_{5,r}\geq h\geq 3$. Because $n\geq T_{1,r}$, by Lemma 2.6 and Remark 2.7 (2), we have, $|\sigma_{R}|=k_{1}-l\geq A_{1}+A_{5,r}+3\geq A_{1}+h+3$. Thus,
$\sigma_{y}=(j_{1},\ldots,j_{l},j_{l+1},j_{l+2},\ldots,j_{l+h-2},j_{0},j_{0},j_{l+h+1}\ldots,j_{k_{1}}),$
$\omega_{y}=(j_{1},\ldots,j_{l},j_{0},j_{0},\hat{j}_{l+3},\hat{j}_{l+4},\ldots,\hat{j}_{l+h},\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{2}}).$
We define $\breve{\omega}:=\sigma_{L}\ast(j_{0},j_{0},\hat{j}_{l+3},\ldots,\hat{j}_{l+h},\hat{j}_{l+h+1},\ldots,\hat{j}_{k_{1}})$. Then we have $F_{\omega}\subset F_{\breve{\omega}}$. Since $k_{1}>l+h$, by (2.6)-(2.8), $(\breve{\omega})_{R}$ is a permutation of $\sigma_{R}$. Note that, by Lemma 2.4 (ii), we have, $|\sigma^{\flat}|\geq|\sigma|-A_{1}\geq A_{5,r}+3>h+2$. Thus, $(\breve{\omega}^{\flat})_{R}$ is also a permutation of $(\sigma^{\flat})_{R}$. It follows that
$b_{\breve{\omega}_{y}^{-}}=b_{\sigma_{y}^{-}}\geq a_{\sigma_{L}}=a_{\breve{\omega}_{L}}>b_{\sigma_{y}}=b_{\breve{\omega}_{y}}.$
$\mathcal{E}_{r}(\breve{\omega}^{\flat})=\mathcal{E}_{r}(\sigma^{\flat})\geq\underline{\eta}_{r}^{n}>\mathcal{E}_{r}(\sigma)=\mathcal{E}_{r}(\breve{\omega}).$
This implies that $\breve{\omega}\in\Lambda_{n,r}$ and $\omega\notin\Lambda_{n,r}$, contradicting the hypothesis. ∎
In the remaining part of this section, we assume that $n\geq T_{2,r}:=T_{1,r}+2A_{2}$. Let $\overline{\sigma}\in\mathcal{B}_{n,r}$ be as defined in (2.3). For every $1\leq h\leq 2A_{2}$ , we fix an integer $i_{l+1+h}\in[1,n_{j_{l+1+h}}]$ and define (cf. Lemma 2.4 (ii))
[…]
Then we have $F_{\overline{\sigma}^{*}}\subset F_{\overline{\sigma}}$. We define
(2.9) $\mathcal{F}_{n,r}:=\{\overline{\sigma}*:\overline{\sigma}\in\mathcal{B}_{n,r}\}.$
Remark 2.15.
(1) From the definitions of $\mathcal{B}_{n,r}$ and $\mathcal{F}_{n,r}$, one can easily see that, ${\rm card}(\mathcal{F}_{n,r})={\rm card}(\mathcal{B}_{n,r})=\varphi_{n,r}$. (2) Let $K_{n,r}:=\bigcup_{\overline{\sigma}^{*}\in\mathcal{F}_{n,r}}F_{\overline{\sigma}^{*}}$. By the definition of $\overline{\sigma}^{*}$, we have, $|\overline{\sigma}^{*}_{L}|-|\sigma_{L}|\leq 2+2A_{2}$. This and Lemma 2.4 (i) yield that
$\mu(K_{n,r})\geq\underline{p}^{2(A_{2}+1)(1+A_{1})}\sum_{\sigma\in\Lambda_{n,r}}\mu(F_{\sigma})\geq\underline{p}^{8A_{1}A_{2}}.$
(3) As we showed in [20, Lemma 4], there exists a positive number $D_{L}$, such that, for every $\alpha_{L}\subset\mathbb{R}^{2}$ with cardinality $L$, the following holds:
$\int_{F_{\overline{\sigma}^{*}}}d(x,\alpha_{L})^{r}d\mu(x)\geq D_{L}\mathcal{E}_{r}(\overline{\sigma}^{*}).$
Lemma 2.16.
For every pair $\overline{\sigma}^{*},\overline{\omega}^{*}$ of distinct words in $\mathcal{F}_{n,r}$, we have
$d(F_{\overline{\sigma}^{*}},F_{\overline{\omega}^{*}})\geq(1+\underline{b}^{-2})^{-1}\underline{b}^{2}\max\{|F_{\overline{\sigma}^{*}}|,|F_{\overline{\omega}^{*}}|\}.$
Proof.
If $(\overline{\sigma})_{y},(\overline{\omega})_{y}$ are incomparable, so are $(\overline{\sigma}^{*})_{y},(\overline{\omega}^{*})_{y}$. We have
$d(F_{\overline{\sigma}^{*}},F_{\overline{\omega}^{*}})$ $\geq$ $d_{v}(F_{\overline{\sigma}^{*}},F_{\overline{\omega}^{*}})\geq(1+\underline{b}^{-2})^{-1}\underline{b}^{2}\max\{|F_{\overline{\sigma}}|,|F_{\overline{\omega}}|\}$
$\geq$ $(1+\underline{b}^{-2})^{-1}\underline{b}^{2}\max\{|F_{\overline{\sigma}^{*}}|,|F_{\overline{\omega}^{*}}|\}.$
If $(\overline{\sigma})_{y},(\overline{\omega})_{y}$ are comparable, then by Lemmas 2.12, 2.14, $\sigma_{L},\omega_{L}$ are incomparable. Thus, from Lemma 2.11, we have
$d(F_{\overline{\sigma}^{*}},F_{\overline{\omega}^{*}})\geq d(F_{\overline{\sigma}},F_{\overline{\omega}})\geq 2^{-\frac{1}{2}}\underline{a}^{2}\max\{|F_{\overline{\sigma}^{*}}|,|F_{\overline{\omega}^{*}}|\}.$
Note that $\underline{a}\leq\underline{b}$. The proof of the lemma is complete. ∎
With the above preparations, we can now establish a characterization for the quantization error by applying [20, Lemma 3].
Proposition 2.17.
We have $e_{\varphi_{n,r},r}^{r}(\mu)\asymp\sum_{\sigma\in\Lambda_{n,r}}\mathcal{E}_{r}(\sigma)$.
Proof.
For every $\sigma\in\Lambda_{n,r}$, let $C_{\sigma}$ be an arbitrary point in $F_{\sigma}$. We have
$e_{\varphi_{n,r},r}^{r}(\mu)\leq\sum_{\sigma\in\Lambda_{n,r}}\mu(F_{\sigma})|F_{\sigma}|^{r}\lesssim\sum_{\sigma\in\Lambda_{n,r}}\mathcal{E}_{r}(\sigma).$
Let $\mathcal{F}_{n,r}$ be as defined in (2.9). For distinct words $\overline{\sigma}^{*},\overline{\omega}^{*}\in\mathcal{F}_{n,r}$, we have
(2.10) $\mathcal{E}_{r}(\overline{\sigma}^{*})\geq\underline{\eta}_{r}^{2(A_{2}+1)}\mathcal{E}_{r}(\sigma)\geq\underline{\eta}_{r}^{2(A_{2}+1)+1}\mathcal{E}_{r}(\omega)\geq\underline{\eta}_{r}^{2A_{2}+3}\mathcal{E}_{r}(\overline{\omega}^{*}).$
By (2.10), Lemma 2.16 and Remark 2.15 (3), the assumptions in Lemma 3 of [20] are fulfilled for the conditional measure $\mu_{n,r}:=\mu(\cdot|K_{n,r})$. It follows that
$e^{r}_{\varphi_{n,r},r}(\mu)\geq\mu(K_{n,r})e^{r}_{\varphi_{n,r},r}(\mu_{n,r})\gtrsim\sum_{\sigma\in\mathcal{F}_{n,r}}\mathcal{E}_{r}(\overline{\sigma}^{*})\gtrsim\sum_{\sigma\in\Lambda_{n,r}}\mathcal{E}_{r}(\sigma).$
This completes the proof of the proposition. ∎
3. Auxiliary coding space and auxiliary measures
3.1. Auxiliary coding space
Let $G,G_{y}$ be endowed with discrete topology and let $G^{\mathbb{N}},G_{y}^{\mathbb{N}}$ be endowed with product topology. Then $G^{\mathbb{N}},G_{y}^{\mathbb{N}}$ are both metrizable. The corresponding product metric is compatible with the product topology on $\Phi_{\infty}$. Thus, $\Phi_{\infty}$ is a compact metric space.
With the next lemma, we show that, if $\sigma,\omega\in\Phi^{*}$ and $\sigma]\cap[\omega]\neq\emptyset$, then we have either $\sigma]\subset[\omega]$, or $\omega]\subset[\sigma]$. The proof of this lemma is different from that for Lemma 2.2, because $\Psi^{*},\Phi^{*}$ are endowed with different partial orders.
Lemma 3.1.
(1) Let $\sigma,\omega\in\Phi^{*}$. Assume that $\sigma_{L}\preceq\omega_{L}$ and $\sigma_{R},\omega_{R}$ are comparable, then we have $\sigma_{R}\preceq\omega_{R}$. (2) For every pair $\sigma,\omega\in\Phi^{*}$, we have either $\sigma]\cap[\omega]=\emptyset$, or $\sigma]\subset[\omega]$, or $\omega]\subset[\sigma]$.
Proof.
(1) Assume that $\sigma_{L}\preceq\omega_{L},\omega_{R}\precneqq\sigma_{R}$. For some $p_{1}\geq 0$ and $p_{2}\geq 1$, we write
$\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\times(j_{l+1},\ldots,j_{k});$
$\omega=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),\ldots,(\hat{i}_{l+p_{1}},\hat{j}_{l+p_{1}}))\times(j_{l+1},\ldots,j_{k-p_{2}}).$
Then by the definition of $\Phi_{l}$, we have, $b_{\sigma_{y}^{-}}\geq a_{\sigma_{L}}>b_{\sigma_{y}}$. If $p_{1}=0$, we have
$b_{\omega_{y}}$ $=$ $\prod_{h=1}^{l}b_{j_{h}}\prod_{h=l+1}^{k-p_{2}}b_{j_{h}}=\prod_{h=1}^{k-p_{2}}b_{j_{h}}\geq\prod_{h=1}^{l}a_{i_{h}j_{h}}=a_{\omega_{L}}.$
This contradicts the fact that $\omega\in]\phi_{l}$. Next, we assume that $p_{1}\geq 1$. Note that $a_{ij}<b_{j}$ for every $(i,j)\in G$. It follows that
$b_{\omega_{y}}$ $=$ $\prod_{h=1}^{l}b_{j_{h}}\prod_{h=l+1}^{l+p_{1}}b_{\hat{j}_{h}}\prod_{h=l+1}^{k-p_{2}}b_{j_{h}}$
$=$ $\prod_{h=1}^{k-p_{2}}b_{j_{h}}\prod_{h=l+1}^{l+p_{1}}b_{\hat{j}_{h}}\geq\prod_{h=1}^{l}a_{i_{h}j_{h}}\prod_{h=l+1}^{l+p_{1}}b_{\hat{j}_{h}}$
$>$ $\prod_{h=1}^{l}a_{i_{h}j_{h}}\prod_{h=l+1}^{l+p_{1}}a_{\hat{j}_{h}\hat{j}_{h}}.$
This contradicts the fact that $\omega\in\Phi^{*}$. It follows that $|\omega_{R}|\geq|\sigma_{R}|$ and $\sigma_{R}\preceq\omega_{R}$.
(2) If both $\sigma_{L},\omega_{L}$, and $\sigma_{R},\omega_{R}$, are comparable, then by (1), we have $\sigma]\subset[\omega]$, or $\omega]\subset[\sigma]$. Otherwise, either $\sigma_{L},\omega_{L}$, or $\sigma_{R},\omega_{R}$, are incomparable, and then we have $\sigma]\cap[\omega]=\emptyset$. ∎
Remark 3.2.
Based on Lemma 3.1, we obtain the following useful facts.
  1. (r1)
    For every pair of distinct words $\sigma,\omega\in\Phi_{l}$, we have that $\sigma]\cap[\omega]=\emptyset$. In fact, if $\sigma_{L}\neq\omega_{L}$, then we certainly have that $\sigma]\cap[\omega]=\emptyset$, since $|\sigma_{L}|=|\omega_{L}|=l$; if $\sigma_{L}=\omega_{L}$, then $\sigma_{R},\omega_{R}$ are incomparable, and we again have $\sigma]\cap[\omega]=\emptyset$.
  2. (r2)
    For every $\tau\in G^{l}$, we define $\Omega(\tau):=\{\omega\in G_{y}^{*}:\tau\times\omega\in\Phi_{l}\}$. Then we have
    $\Omega(\tau):=\{\omega\in G_{y}^{*}:b_{\omega^{-}}\geq\frac{a_{\tau}}{b_{\tau_{y}}}>b_{\omega}\}.$
    Hence, $G_{y}^{\mathbb{N}}$ is the disjoint union of the sets $\omega],\omega\in\Omega(\tau)$. Therefore,
    $\bigcup_{\sigma\in\Phi_{l}}[\sigma]=\bigcup_{\tau\in G^{l}}\bigcup_{\omega\in\Omega(\tau)}[\tau\times\omega]=\Phi_{\infty}.$
Remark 3.3.
The following facts have been noted in [30] and will also be useful for the proof of the main theorem.
  1. (r3)
    It can happen that for some $\sigma=\sigma_{L}\ast\sigma_{R},\omega=\omega_{L}\ast\omega_{R}\in\Psi^{*}$ with $F_{\sigma}^{\circ}\cap F_{\omega}^{\circ}=\emptyset$, but $\mathcal{L}(\sigma)]\supset[\mathcal{L}(\omega)]$. This can be seen by considering
    $\sigma_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l})),\;\sigma_{R}=(j_{l+1},\ldots,j_{k});\;j_{l+1}\neq\hat{j}_{l+1};$
    $\omega_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(i_{l+1},\hat{j}_{l+1})),\sigma_{R}=(j_{l+1},\ldots,j_{k},\ldots,j_{k+p}).$
  2. (r4)
    It can happen that for some $\sigma=\sigma_{L}\times\sigma_{R},\omega=\omega_{L}\times\omega_{R}\in\Phi^{*}$, $\sigma]\cap[\omega]=\emptyset$, but $F_{\mathcal{L}^{-1}(\omega)}\subset F_{\mathcal{L}^{-1}(\sigma)}$. This can be seen by considering
    $\sigma_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l})),\;\sigma_{R}=(j_{l+1},\ldots,j_{k});\;j_{l+1}\neq j_{l+2};$
    $\omega_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(i_{l+1},j_{l+1})),\;\sigma_{R}=(j_{l+2},\ldots,j_{k},\ldots,j_{k+p}).$
3.2. Auxiliary measures
For $h\geq 1,l\geq 1$ and $\sigma\in\Phi_{l}$, we define
$\mathcal{E}_{r}(\sigma):=\mathcal{E}_{r}(\mathcal{L}^{-1}(\sigma)),\;\Lambda_{h}(\sigma):=\{\rho\in\Phi_{l+h}:\sigma\preceq\rho\};$
$I_{h,r}(t):=\sum_{\omega\in\Phi_{h}}\mathcal{E}_{r}(\omega)^{t},\;t>0.$
In order to construct an auxiliary measure on $\Phi_{\infty}$, we need to prove
(3.1) $\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\asymp\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$
We divide the proof of (3.1) into the subsequent three lemmas.
Lemma 3.4.
Let $f_{1}(t):=\min\{1,(\sum_{j\in G_{y}}q_{j}^{t})^{-1}\}$. For $\sigma\in\Phi^{*}$ and $h\geq 1$, we have
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\geq f_{1}(t)\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$
Proof.
[…]
(3.2) $\prod_{p=1}^{k-1}b_{j_{p}}\geq\prod_{p=1}^{l}a_{i_{p}j_{p}}>\prod_{p=1}^{k}b_{j_{p}};\;\;\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\geq\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>\prod_{p=1}^{\hat{k}}b_{\hat{j}_{p}}.$
It follows that
$\prod_{p=1}^{k-1}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\geq\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}}b_{j_{p}}.$
We need to distinguish between the following two cases:
  1. (1)
    $\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{j_{p}}\geq\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}$. In this case, we define
    $\rho(\omega):=(\sigma_{L}\ast\omega_{L})\times(\sigma_{R}\ast\omega_{R}).$
    We have that $\rho(\omega)\in\Lambda_{h}(\sigma)$. We denote by $\Lambda_{h,1}(\sigma)$ the of set such words $\rho(\omega)$ and denote the set of the words $\omega$ by $\Phi_{h,1}$.
  2. (2)
    $\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}<\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}$. We define
    $\rho(\omega):=(\sigma_{L}\ast\omega_{L})\times(j_{l+1},\ldots,j_{k-1},\hat{j}_{h+1},\ldots,\hat{j}_{\hat{k}-1},j_{k}).$
    Then $\rho(\omega)\in\Lambda_{h}(\sigma)$. We denote the set of such words $\rho(\omega)$ by $\Lambda_{h,2}(\sigma)$ and denote the set of the words $\omega$ by $\Phi_{h,2}$.
Let $\tau\in G_{y}^{*}$. If $|\tau|=1$, we define $\tau^{\sharp}:=\theta$; otherwise, we denote by $\tau^{\sharp}$ the word that is obtained by deleting the first letter of $\tau$. We observe the following facts.
  1. (i)
    We have, $\Phi_{h,1}\cap\Phi_{h,2}=\emptyset,\Lambda_{h,2}(\sigma)\cap\Lambda_{h,2}(\sigma)=\emptyset$. The first equality is clear, so we need only to show the second. In fact, a word $\rho(\omega)\in\Lambda_{h,1}(\sigma)$ can not be obtained by any $\tau\in\Phi_{h,2}$ and vice versa. Otherwise, we would have $\omega_{L}=\tau_{L},\tau_{R}|_{1}=j_{k}$ and $\omega_{R}=(\tau_{R}^{-})^{\sharp}\ast j_{k}$. This leads to
    $b_{\omega_{y}}=b_{(\omega_{L})_{y}}b_{\omega_{R}}=b_{(\tau_{L})_{y}}b_{\omega_{R}}=b_{(\tau_{L})_{y}}b_{(\tau_{R}^{-})^{\sharp}\ast j_{k}}=b_{\tau^{-}}\geq a_{\tau_{L}}=a_{\omega_{L}},$
    contradicting the fact that $\omega\in\Phi_{h}$.
  2. (ii)
    For different words $\omega,\tau\in\Phi_{h,1}$, we have $\rho(\omega)\neq\rho(\tau)$.
  3. (iii)
    There exist at most $m$ words in $\Phi_{h,2}$ that determine the same word in $\Lambda_{h,2}(\sigma)$, because of the absence of $\hat{j}_{\hat{k}}$ in $\rho(\omega)$. Fix an $\omega\in\Phi_{h,2}$. For $j\in G_{y}$, let $\hat{\omega}^{(j)}:=\omega_{L}\times(\omega_{R}^{-}\ast j)$. Whenever $\hat{\omega}^{(j)}\in\Phi_{h}$, we have, $\rho(\hat{\omega}^{(j)})=\rho(\omega)$ and $\hat{\omega}^{(j)}\in\Phi_{h,2}$. Hence, we obtain that (cf. Remark 3.5 below)
    $\rho^{-1}(\rho(\omega))=\{\hat{\omega}^{(j)}:j\in G_{y}\}\cap\Phi_{h}.$
    We write $\langle\omega\rangle:=\{\hat{\omega}:\rho(\hat{\omega})=\rho(\omega)\}$. For every $\omega\in\Phi_{h}(2)$, we take an arbitrary word of $\langle\omega\rangle$ and denote the set of all such words by $\Phi^{\flat}_{h}(2)$.
Using the above facts, we deduce that
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}$ $\geq$ $\sum_{\rho\in\Lambda_{h,1}(\sigma)}\mathcal{E}_{r}(\rho)^{t}+\sum_{\rho\in\Lambda_{h,2}(\sigma)}\mathcal{E}_{r}(\rho)^{t}$
$=$ $\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi_{h}(1)}\mathcal{E}_{r}(\omega)^{t}+\sum_{\omega\in\Phi^{\flat}_{h}(2)}\mathcal{E}_{r}(\rho(\omega))^{t}$
$=$ $\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi_{h}(1)}\mathcal{E}_{r}(\omega)^{t}+\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi^{\flat}_{h}(2)}(p_{\omega_{L}}q_{\omega_{R}^{-}}a_{\omega_{L}}^{r})^{t}$
$\geq$ $\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi_{h}(1)}\mathcal{E}_{r}(\omega)^{t}+\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi_{h}(2)}\mathcal{E}_{r}(\omega)^{t}\big{(}\sum_{j\in G_{y}}q_{j}^{t}\big{)}^{-1}$
$\geq$ $f_{1}(t)\mathcal{E}_{r}(\sigma)^{t}\sum_{\omega\in\Phi_{h}}\mathcal{E}_{r}(\omega)^{t}.$
This completes the proof of the lemma. ∎
Remark 3.5.
Assume that $\omega\in\Phi_{h,2}$ and for some $j\in G_{y}$, $\hat{\omega}^{(j)}\notin\Phi_{h}$. Then there exists some $p\geq 1$ such that $\hat{\omega}^{(j)+}:=\omega_{L}\times(\omega_{R}^{-}\ast(j,\widetilde{j}_{1},\ldots,\widetilde{j}_{p}))\in\Phi_{h,2}$. To see this, let $\sigma,\omega$ be the same as in Lemma 3.4. Then (3.2) holds. Since $\hat{\omega}^{(j)}\notin\Phi_{h}$, we have, $\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}b_{j}>\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}$. Thus, there exists some integer $p\geq 1$, such that
$\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\cdot b_{j\ast\widetilde{j}_{1},\ldots,\widetilde{j}_{p-1}}\geq\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\cdot b_{j\ast\widetilde{j}_{1},\ldots,\widetilde{j}_{p}}.$
Since $\omega\in\Phi_{h,2}$, we deduce that
$\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\cdot b_{j\ast\hat{j}_{1}\cdots\ast\hat{j}_{p-1}}<\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}<\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}.$
This implies that $\omega_{L}\times(\omega_{R}^{-}\ast(j,\widetilde{j}_{1},\ldots,\widetilde{j}_{p}))\in\Phi_{h,2}$. Note that
$\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}<\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}\leq\prod_{p=1}^{k-1}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\cdot b_{j}.$
We clearly have $b_{j}>b_{\hat{j}_{\hat{k}}}$. We have
$\rho(\hat{\omega}^{(j)+})=(\sigma_{L}\ast\omega_{L})\times(\sigma_{R}\ast\omega_{R}^{-}\ast(j,\widetilde{j}_{1},\ldots,\widetilde{j}_{p-1},j_{k}).$
Analogously to Lemma 2.4, we have
Lemma 3.6.
Let $A_{6}:=[A_{4}^{-1}]$ and $\sigma\in\Phi^{*}$. Then (1) for every $h>A_{6}$ and $\rho\in\Lambda_{h}(\sigma)$, we have $|\rho_{R}|\geq|\sigma_{R}|+1$; (2) for every $\rho\in\Lambda_{1}(\sigma)$, we have $|\rho_{R}|-|\sigma_{R}|\leq A_{1}$.
Proof.
Let $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\times(j_{l+1},\ldots,j_{k})\in\Phi_{l}$. Assume that $h>A_{6}$ and $\rho\in\Lambda_{h}(\sigma)$. We write $\rho_{L}:=\sigma_{L}\ast((\hat{i}_{1},\hat{j}_{1}),\ldots,(\hat{i}_{h},\hat{j}_{h}))$. Then we have
$\prod_{p=1}^{l+h}a_{i_{h}j_{h}}$ $=$ $\prod_{p=1}^{l}a_{i_{h}j_{h}}\prod_{p=1}^{h}a_{\hat{i}_{h}\hat{j}_{h}}\leq\prod_{p=1}^{k-1}b_{j_{h}}\prod_{p=1}^{h}a_{i_{h}j_{h}}$
$\leq$ $\underline{b}^{-1}\prod_{p=1}^{k}b_{j_{h}}\prod_{p=1}^{h}a_{\hat{i}_{h}\hat{j}_{h}}=\underline{b}^{-1}\prod_{p=1}^{k}b_{j_{h}}\prod_{p=1}^{h}b_{\hat{j}_{p}}\prod_{p=1}^{h}\frac{a_{\hat{i}_{h}\hat{j}_{h}}}{b_{\hat{j}_{p}}}$
$\leq$ $\underline{b}^{-1}A_{3}^{h}\prod_{p=1}^{k}b_{j_{h}}\prod_{p=1}^{h}b_{\hat{j}_{p}}$
$<$ $\prod_{p=1}^{k}b_{j_{h}}\prod_{p=1}^{h}b_{\hat{j}_{p}}$
It follows that $|\rho_{R}|>k-l$ and (1) follows. (2) can be proved similarly. ∎
Remark 3.7.
For $\sigma\in\Phi^{*}$ and $h\geq 1$, we have ${\rm card}(\Lambda_{h}(\sigma))\leq N^{h}m^{hA_{1}}$, where $N:={\rm card}(G)$. Therefore,
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\leq N^{h}m^{hA_{1}}\overline{\eta}_{r}^{ht}\mathcal{E}_{r}(\sigma)^{t}.$
Let $\xi(t):=\max\limits_{1\leq h\leq A_{6}}(I_{h,r}(t)^{-1}N^{h}m^{hA_{1}}\overline{\eta}_{r}^{ht})$. Then for every $1\leq h\leq A_{6}$, we have
$\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\leq\xi(t)\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$
Lemma 3.8.
Let $f_{2}(t):=\max\{\xi(t),\underline{q}^{-t}\}$. For every $\sigma\in\Phi_{l}$ and $h\geq 1$, we have
(3.3) $\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\leq f_{2}(t)\mathcal{E}_{r}(\sigma)^{t}I_{h,r}(t).$
Proof.
Let $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\times(j_{l+1},\ldots,j_{k})\in\Phi_{l}$. By Remark 3.7, it suffices to prove (3.3) for $h\geq A_{6}$. Let $\rho$ be an arbitrary word in $\Lambda_{h}(\sigma)$. We write
$\rho_{L}=((i_{1},j_{1}),\ldots,(i_{l},j_{l}),(\hat{i}_{1},\hat{j}_{1}),\ldots,(\hat{i}_{h},\hat{j}_{h})),$
[…]
By the definition of $\Phi^{*}$, we have
$\prod_{p=1}^{k-1}b_{j_{p}}\geq\prod_{p=1}^{l}a_{i_{p}j_{p}}>\prod_{p=1}^{k}b_{j_{p}};$
(3.4) $\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}\geq\prod_{p=1}^{l}a_{i_{p}j_{p}}\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>\prod_{p=1}^{k}b_{j_{p}}\prod_{p=1}^{\hat{k}}b_{\hat{j}_{p}}.$
If $\hat{k}=1$, we replace $\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}$ in (3.4) with $1$. As a consequence, we obtain
$\prod_{p=1}^{\hat{k}-1}b_{\hat{j}_{p}}>\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>b_{j_{k}}\prod_{p=1}^{\hat{k}}b_{j_{p}}.$
We distinguish between the following two cases.
  1. (i)
    $\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}>\prod_{p=1}^{\hat{k}}b_{j_{p}}$. In this case, we define
[…]
    Then $\omega(\rho)\in\Phi_{h}$. Let $\Phi_{h,1}$ denote the set of all such words $\omega(\rho)$ and let $\Lambda_{h,1}$ denote the set of the words $\rho$ in this case.
  2. (ii)
    $\prod_{p=1}^{h}a_{\hat{i}_{p}\hat{j}_{p}}\leq\prod_{p=1}^{\hat{k}}b_{j_{p}}$. In this case, we define
    $\omega(\rho):=((\hat{i}_{1},\hat{j}_{1}),\ldots,(\hat{i}_{h},\hat{j}_{h}))\times(\hat{j}_{h+1},\ldots,\hat{j}_{\hat{k}},j_{k}).$
    One can see that $\omega(\rho)\in\Phi_{h}$. Let $\Phi_{h,2}$ denote the set of all such words $\omega$ and let $\Lambda_{h,2}$ the set of the words $\rho$ in this case.
We clearly have that $\Lambda_{h,1}\cap\Lambda_{h,2}=\emptyset$. Also, we have $\Phi_{h,1}\cap\Phi_{h,2}=\emptyset$. Otherwise, there would exist some $\rho^{(1)}\in\Lambda_{h,1}$ and $\rho^{(2)}\in\Lambda_{h,2}$ such that $\omega:=\omega(\rho^{(1)})=\omega(\rho^{(2)})=:\tau$. Then we have, $\omega_{L}=\tau_{L}$ and $\tau_{R}=\omega_{R}^{-}$ contradicting (3.4). It follows that
$I_{h,r}(t)$ $\geq$ $\sum_{\omega\in\Phi_{h,1}}\mathcal{E}_{r}(\omega)^{t}+\sum_{\omega\in\Phi_{h,2}}\mathcal{E}_{r}(\omega)^{t}$
$\geq$ $\sum_{\rho\in\Lambda_{h,1}}\frac{\mathcal{E}_{r}(\rho)^{t}}{\mathcal{E}_{r}(\sigma)^{t}}+\underline{q}^{t}\sum_{\rho\in\Lambda_{h,2}}\frac{\mathcal{E}_{r}(\rho)^{t}}{\mathcal{E}_{r}(\sigma)^{t}}$
$\geq$ $\underline{q}^{t}\mathcal{E}_{r}(\sigma)^{-t}\sum_{\rho\in\Lambda_{h}(\sigma)}\mathcal{E}_{r}(\rho)^{t}.$
This completes the proof of the lemma. ∎
Combining Lemmas 3.4, 3.8, we obtain the next lemma, which will be used in the construction of auxiliary measures.
Lemma 3.9.
For $k,p\geq 1$, we have
$f_{1}(t)I_{k,r}(t)I_{p,r}(t)\leq I_{k+p,r}(t)\leq f_{2}(t)I_{k,r}(t)I_{p,r}(t).$
Proof.
Using Lemmas 3.4, 3.8, we have
$I_{k+p,r}(t)=\sum_{\sigma\in\Phi_{k}}\sum_{\rho\in\Lambda_{p}(\sigma)}\mathcal{E}_{r}(\rho)^{t}\geq f_{1}(t)I_{k,r}(t)I_{p,r}(t).$
The second inequality can be obtained similarly. ∎
As a consequence of Lemma 3.9, we can obtain the following standard result.
Proposition 3.10.
(1) For every $t>0$, $\lim_{k\to\infty}\frac{1}{k}\log I_{k,r}(t)=:g(t)$ exists. (2) There exists a unique $s_{r}>0$ such that for $t_{r}=\frac{s_{r}}{s_{r}+r}$, we have $g(t_{r})=0$. (3) For $C(t):=f_{1}(t)^{-1}f_{2}(t)$ and $k,p\geq 1$, we have,
$C(t)^{-1}I_{k}(t_{r})\leq I_{p}(t_{r})\leq C(t)I_{k}(t_{r}).$
Proof.
This can be proved by using Lemma 3.9, the fact that $\dim_{B}E>0$ and [5, Corollary 1.2] and along the line of [5, Lemma 5.2]. ∎
With the help of Lemmas 3.4-3.9, we are now able to construct an auxiliary probability measure on $\Phi_{\infty}$ by applying Prohorov’s theorem. Recall that a family $\mathcal{\pi}$ of probability measures on a metric space $X$, is said to be tight if for every $\epsilon\in(0,1)$, there exists some compact subset $K$ of $X$ such that $\inf_{\nu\in\mathcal{\pi}}\nu(K)\geq 1-\epsilon$. In particular, if $X$ is a compact metric space, then every family $\mathcal{\pi}$ of probability measures on $X$ is tight.
Lemma 3.11.
There exists a Borel probability measure $\lambda$ on $\Phi_{\infty}$ such that, for every $\sigma\in\Phi^{*}$, we have $\lambda([\sigma])\asymp\mathcal{E}_{r}(\sigma)^{t_{r}}$.
Proof.
For every $k\geq 1$ and $\sigma\in\Phi_{k}$, let $C_{\sigma}$ be an arbitrary point in $\sigma]$ and let $\delta_{C_{\sigma}}$ denote the Dirac measure at the point $C_{\sigma}$. We define
$\lambda_{k}=\frac{1}{I_{k,r}(t_{r})}\sum_{\sigma\in\Phi_{k}}\mathcal{E}_{r}(\sigma)^{t_{r}}\delta_{C_{\sigma}}.$
Then $(\lambda_{k})_{k=1}^{\infty}$ is a sequence of probability measures on $\Phi_{\infty}$. Note that $\Phi_{\infty}$ is a compact metric space; so $(\lambda_{k})_{k=1}^{\infty}$ is tight. According to Prohorov’s Theorem (cf. [3, Theorem 5.1]), there exists a subsequence $(\lambda_{k_{i}})_{i=1}^{\infty}$ and a Probability measure $\lambda$ on $\Phi_{\infty}$ such that $\lambda_{k_{i}}$ converges weakly to $\lambda$. Let $k\geq 1$ and $\sigma\in\Phi_{k}$ be given. For every $p>k$, using Lemmas 3.4-3.9, we deduce
$\lambda_{p}([\sigma])$ $=$ $\sum_{\rho\in\Lambda_{p-k}(\sigma)}\lambda_{p}([\rho])=\frac{1}{I_{p,r}(t_{r})}\sum_{\rho\in\Lambda_{p-k}(\sigma)}\mathcal{E}_{r}(\rho)^{t_{r}}$
$\asymp$ $\frac{1}{I_{p,r}(t_{r})}\mathcal{E}_{r}(\sigma)^{t_{r}}I_{p-k,r}(t_{r})\asymp\mathcal{E}_{r}(\sigma)^{t_{r}}.$
Note that $\sigma_{L}]$ ($\sigma_{R}]$) is both open and closed in $G^{\mathbb{N}}$ ($G_{y}^{\mathbb{N}}$). Thus $\sigma]$ is clopen in $\Phi_{\infty}$ (for closeness, it suffices to note that its complement is open). Because $\Phi_{\infty}$ is compact, we deduce that $\sigma]$ is also compact. It follows that $\lambda([\sigma])\asymp\mathcal{E}_{r}(\sigma)^{t_{r}}$. ∎
4. Proof of Theorem 1.2
By [31, Lemma 3.4], for the proof of Theorem 1.2, it is sufficient to show
(4.1) $0<\liminf_{n\to\infty}\sum_{\sigma\in\Lambda_{n,r}}\mathcal{E}_{r}(\sigma)^{t_{r}}\leq\limsup_{n\to\infty}\sum_{\sigma\in\Lambda_{n,r}}\mathcal{E}_{r}(\sigma)^{t_{r}}<\infty.$
Next, we give the proof for the first inequality of (4.1), in an analogous manner to that for [30, Proposition 3.2].
Lemma 4.1.
We have $\sum_{\sigma\in\Lambda_{n,r}}\mathcal{E}_{r}(\sigma)^{t_{r}}\lesssim 1$.
Proof.
For every $n\geq 1$, let $\Lambda_{n,r}$ be as defined in (2.2). We define
$\hat{\Lambda}_{n,r}:=\{\hat{\sigma}=\mathcal{L}(\sigma):\sigma\in\Lambda_{n,r}\}.$
By Lemma 3.1, for $\hat{\sigma},\hat{\omega}\in\hat{\Lambda}_{n,r}$, the sets $\hat{\sigma}],[\hat{\omega}$], are either disjoint, or one is contained in the other. Also, by Remark 2.7 (1), we have
$\underline{\eta}_{r}\mathcal{E}_{r}(\hat{\omega})\leq\mathcal{E}_{r}(\hat{\sigma})\leq\underline{\eta}_{r}^{-1}\mathcal{E}_{r}(\hat{\omega}),\;\hat{\sigma},\hat{\omega}\in\hat{\Lambda}_{n,r}.$
If $\hat{\omega}]\subset[\hat{\sigma}$], then as we did in [30, Lemma 3.1], one can see that, there exists some constant $H_{1,r}\geq 1$ such that $\big{|}|\hat{\sigma}_{L}|-|\hat{\omega}_{L}|\big{|}\leq H_{1,r}$. This allows us to select a subset $\hat{\Lambda}_{n,r}^{\flat}$ of $\hat{\Lambda}_{n,r}$ such that, the sets $\hat{\sigma}],\hat{\sigma}\in\hat{\Lambda}_{n,r}^{\flat}$, are pairwise disjoint, and
$\hat{\Lambda}_{n,r}=\bigcup_{\hat{\sigma}\in\hat{\Lambda}_{n,r}^{\flat}}\Gamma(\hat{\sigma}),\;{\rm with}\;\;\Gamma(\hat{\sigma})=\{\mathcal{L}(\omega):\hat{\sigma}\preceq\mathcal{L}(\omega),\omega\in\Lambda_{n,r}\}.$
Combining the preceding equality with Lemma 3.11, we obtain
$\sum_{\hat{\sigma}\in\hat{\Lambda}_{n,r}}\mathcal{E}_{r}(\hat{\sigma})^{t_{r}}\asymp\sum_{\sigma\in\hat{\Lambda}_{n,r}}\lambda([\hat{\sigma}])\leq H_{1,r}\sum_{\sigma\in\hat{\Lambda}_{n,r}^{\flat}}\lambda([\hat{\sigma}])\leq H_{1,r}.$
This completes the proof for the lemma. ∎
In the following, we are going to prove the last inequality in (4.1). We need to define the predecessors for $\sigma\in\Phi^{*}$. For $\sigma\in\Phi_{1}$, we define $\sigma^{-}:=\theta$. Let $l\geq 2$ and $\sigma=((i_{1},j_{1}),\ldots,(i_{l},j_{l}))\times(j_{l+1},\ldots,j_{k})\in\Phi_{l}$, we have, $b_{\sigma_{y}^{-}}\geq a_{\sigma_{L}}>b_{\sigma_{y}}$. Hence,
$\prod_{h=1}^{l-1}a_{i_{h}j_{h}}>a_{i_{l}j_{l}}^{-1}\prod_{h=1}^{k}b_{j_{h}}=\frac{b_{j_{l}}}{a_{i_{l}j_{l}}}\prod_{h=1}^{l-1}b_{j_{h}}\prod_{h=l+1}^{k}b_{j_{h}}>\prod_{h=1}^{l-1}b_{j_{h}}\prod_{h=l+1}^{k}b_{j_{h}}.$
There exists a unique integer $h\geq 0$, such that the following inequalities hold:
$\prod_{h=1}^{l-1}b_{j_{h}}\prod_{h=l+1}^{k-h-1}b_{j_{h}}\geq\prod_{h=1}^{l-1}a_{i_{h}j_{h}}>\prod_{h=1}^{l-1}b_{j_{h}}\prod_{h=l+1}^{k-h}b_{j_{h}}.$
For this integer $h$, we define $\sigma^{-}:=\sigma_{L}^{-}\times(j_{l+1},\ldots,j_{k-h})$. By Lemma 3.6, we obtain
(4.2) $\underline{\eta}_{r}\mathcal{E}_{r}(\sigma^{-})\leq\mathcal{E}_{r}(\sigma)\leq\overline{p}\;\overline{a}^{r}\mathcal{E}_{r}(\sigma)<\overline{\eta}_{r}\mathcal{E}_{r}(\sigma^{-}).$
Let $\mathcal{S}_{n,r}:=\{\sigma\in\Psi^{*}:\underline{\eta}_{r}^{n+1}\leq\mathcal{E}_{r}(\sigma)<\underline{\eta}_{r}^{n}\}$. We define
(4.3) $G_{1}(\sigma):=\{\omega\in\mathcal{S}_{n,r}:\sigma\preceq\omega\},\;\sigma\in\mathcal{S}_{n,r}.$
Remark 4.2.
Let $M_{r}:=[\frac{\log\underline{\eta}_{r}}{\log\overline{\eta}_{r}}]$. For every $\omega\in G_{1}(\sigma)$, we have, $|\omega_{L}|-|\sigma_{L}|\leq M_{r}$. In fact, assume that $|\omega_{L}|-|\sigma_{L}|>M_{r}$. Then we have
$\mathcal{E}_{r}(\omega)\leq\mathcal{E}_{r}(\sigma)\cdot\overline{\eta}_{r}^{M_{r}}<\underline{\eta}_{r}^{n+1}.$
This contradicts the definition of $G_{1}(\sigma)$. If follows that $G_{1}(\sigma)\subset\bigcup_{h=1}^{M_{r}}\Gamma_{h}(\sigma)$, where
(4.4) $\Gamma_{h}(\sigma):=\{\rho\in\Psi_{l+h}:\sigma\preceq\rho\}.$
The following lemma is an analogue of [32, Lemma 4.1].
Lemma 4.3.
There exists a constant $H_{2,r}>0$ such that, for every $\sigma\in\mathcal{S}_{n,r}$,
$\sum_{\omega\in G_{1}(\sigma)}\mathcal{E}_{r}(\omega)^{t_{r}}\leq H_{2,r}\mathcal{E}_{r}(\sigma)^{t_{r}}.$
Proof.
By Lemma 2.4 (i) and a coarse estimate, for every $h\geq 1$, we have that ${\rm card}(\Gamma_{h}(\sigma))\leq N^{h}m^{hA_{1}}$. This, along with (2.1), yields
$\sum_{\omega\in G_{1}(\sigma)}\mathcal{E}_{r}(\omega)^{t_{r}}\leq\sum_{h=1}^{M_{r}}\sum_{\omega\in\Gamma_{h}(\sigma)}\mathcal{E}_{r}(\omega)^{t_{r}}\leq\sum_{h=1}^{M_{r}}{\rm card}(\Gamma_{h}(\sigma))\overline{\eta}_{r}^{ht_{r}}\mathcal{E}_{r}(\sigma)^{t_{r}}.$
It is sufficient to define $H_{2,r}:=\sum_{h=1}^{M_{r}}N^{h}m^{hA_{1}}\overline{\eta}_{r}^{ht_{r}}$. ∎
Lemma 4.4.
We have $\sum_{\sigma\in\Lambda_{n,r}}\mathcal{E}_{r}(\sigma)^{t_{r}}\gtrsim 1$.
Proof.
For every $n\geq 1$, we define
$\Gamma_{n,r}:=\{\sigma\in\Phi^{*}:\mathcal{E}_{r}(\sigma^{-})\geq\underline{\eta}_{r}^{n}>\mathcal{E}_{r}(\sigma)\}.$
By Lemma 3.1 and Remark 3.2 (r2), for every pair $\sigma,\omega\in\Gamma_{n,r}$, we have
$\sigma]\cap[\omega]=\emptyset,\;\bigcup_{\sigma\in\Gamma_{n,r}}[\sigma]=\Phi_{\infty}.$
From this and Lemma 3.11, it follows that
(4.5) $\sum_{\sigma\in\Gamma_{n,r}}\mathcal{E}_{r}(\sigma)^{t_{r}}\asymp\sum_{\sigma\in\Gamma_{n,r}}\lambda([\sigma])=1.$
Now we connect the words in $\Gamma_{n,r}$ with the approximate squares. We define
$\widetilde{\Gamma}_{n,r}:=\{\widetilde{\sigma}:=\mathcal{L}^{-1}(\sigma)=\sigma_{L}\ast\sigma_{R}:\sigma\in\Gamma_{n,r}\}.$
From (4.2) and the definition of $\Gamma_{n,r}$, we know that, $\widetilde{\Gamma}_{n,r}\subset\mathcal{S}_{n,r}$; and every $\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}$ corresponds to an approximate square. In view of Remark 3.2 (r2), we define
$G(\widetilde{\sigma})=\{\widetilde{\omega}\in\widetilde{\Gamma}_{n,r}:F_{\widetilde{\omega}}\subset F_{\widetilde{\sigma}}\},\;\;\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}.$
[…]
Using this, Lemma 4.3 and (4.5), we deduce
(4.6) $\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}\geq H_{2,r}^{-1}\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}=H_{2,r}^{-1}\sum_{\sigma\in\Gamma_{n,r}}\mathcal{E}_{r}(\sigma)^{t_{r}}\asymp 1.$
Next, we compare the words in $\widetilde{\Gamma}_{n,r}^{\flat}$ with those in $\Lambda_{n,r}$. We define
$\Lambda_{n,r}^{\flat}:=\{\sigma\in\Lambda_{n,r}:F^{\circ}_{\sigma}\cap\bigg{(}\bigcup_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}}F_{\widetilde{\sigma}}^{\circ}\bigg{)}\neq\emptyset\}$.
We need to divide $\Lambda_{n,r}^{\flat}$ and $\widetilde{\Gamma}_{n,r}^{\flat}$ into two subsets:
$\Lambda_{n,r}^{\flat}(1):=\{\sigma\in\Lambda_{n,r}:F_{\sigma}\subseteq F_{\widetilde{\sigma}}\;{\rm for\;some}\;\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}\},$
$\Lambda_{n,r}^{\flat}(2):=\{\sigma\in\Lambda_{n,r}:F_{\sigma}\supsetneq F_{\widetilde{\sigma}}\;{\rm for\;some}\;\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}\},$
$\widetilde{\Gamma}_{n,r}^{\flat}(1):=\{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}:F_{\sigma}\subseteq F_{\widetilde{\sigma}}\;{\rm for\;some}\;\sigma\in\Lambda_{n,r}\},$
$\widetilde{\Gamma}_{n,r}^{\flat}(2):=\{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}:F_{\sigma}\supsetneq F_{\widetilde{\sigma}}\;{\rm for\;some}\;\sigma\in\Lambda_{n,r}\}.$
By Lemma 3.1, $\Lambda_{n,r}^{\flat}$ is the disjoint union of $\Lambda_{n,r}^{\flat}(1)$ and $\Lambda_{n,r}^{\flat}(2)$. We further define
$S(\sigma):=\{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}:F_{\sigma}\supsetneq F_{\widetilde{\sigma}}\};\;\sigma\in\Lambda_{n,r}^{\flat}(2);$
$T(\widetilde{\sigma}):=\{\sigma\in\Lambda_{n,r}:F_{\sigma}\subseteq F_{\widetilde{\sigma}}\},\;\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}(1).$
[…]
$A_{1,r}:=\sum_{\sigma\in\Lambda_{n,r}^{\flat}(1)}\mathcal{E}_{r}(\sigma)^{t_{r}};\;A_{2,r}:=\sum_{\sigma\in\Lambda_{n,r}^{\flat}(2)}\mathcal{E}_{r}(\sigma)^{t_{r}}.$
In the following, we estimate $A_{1,r}$ and $A_{2,r}$ separately.
[…]
$\sum_{\sigma\in T(\widetilde{\sigma})}\mathcal{E}_{r}(\sigma)^{t_{r}}=\sum_{\sigma\in T(\widetilde{\sigma})}(\mu(F_{\sigma})a_{\sigma_{L}^{r}})^{t_{r}}\geq\underline{a}^{\frac{M_{r}rs_{r}}{s_{r}+r}}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}.$
Let $H_{3,r}:=\underline{a}^{\frac{M_{r}rs_{r}}{s_{r}+r}}$. It follows that
(4.7) $A_{1,r}=\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}(1)}\sum_{\sigma\in T(\widetilde{\sigma})}\mathcal{E}_{r}(\sigma)^{t_{r}}\geq H_{3,r}\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}(1)}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}.$
For every $\sigma\in\Lambda_{n,r}^{\flat}(2)$, we have, $S(\sigma)\subseteq G_{1}(\sigma)$. Thus, by Lemma 4.3, we deduce
(4.8) $A_{2,r}\geq\sum_{\sigma\in\Lambda_{n,r}^{\flat}(2)}H_{2,r}^{-1}\sum_{\widetilde{\sigma}\in S(\sigma)}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}=H_{2,r}^{-1}\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}(2)}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}$
Let $H_{4,r}:=\min(H_{2,r}^{-1},H_{3,r})$. Combining (4.6)-(4.8), we obtain
$\sum_{\sigma\in\Lambda_{n,r}}\mathcal{E}_{r}(\sigma)^{t_{r}}\geq A_{1,r}+A_{2,r}\geq H_{4,r}\sum_{\widetilde{\sigma}\in\widetilde{\Gamma}_{n,r}^{\flat}}\mathcal{E}_{r}(\widetilde{\sigma})^{t_{r}}\gtrsim 1.$
This completes the proof of the lemma. ∎
Proof of Theorem 1.2 This is an easy consequence of Remark 2.7, Lemmas 2.17, 4.1, 4.4 and [31, Lemma 3.4].